### v13

# Accurate Low Rank Approximation at a Low Computational Cost

Abstract
Given an algorithm $\mathbb{A}$ for computing a crude Low Rank Approximation (LRA) of a matrix having fast decaying spectra of singular values, we greatly decreased the spectral norm of its output error matrix at sublinear additional computational cost. Furthermore, for a variety of synthetic and real world inputs we tested numerically variants where our entire algorithms run at sublinear cost, and in most cases we still obtained accurate or even near-optimal LRAs. Towards enhancing our progress to a larger class of matrices, we nontrivially extended to LRA the popular technique of iterative refinement.
Key Words:
low rank approximation, near-optimal algorithms, sublinear cost algorithms, singular values, gaps in the spectrum, iterative refinement
2020 Math. Subject Classification:
65F55, 65N75, 65Y20, 68Q25, 68W20
1 LRA of a matrix: State of the Art and our Novelties
1.1 The State of the Art
Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra. [^1] The terms “low”, “large”, “small”, “close”, “near” etc are used in context in our introductory comments. Next we outline its State of the Art, referring for broader expositions to foremost papers [20, 54].
1.1.1. Some assumptions. For simplicity we seek LRA just in the field $\mathbb{R}$ of real numbers but can readily extend our studies to the complex field $\mathbb{C}$. We assume that an $m\times n$ input matrix $M$ is a black box matrix, given by an oracle (black box subroutine) for computing its product with a vector rather than by the entries. This is a natural model, e.g., for the study of matrices of immense sizes coming from Big Data, for which the access to all matrix entries is too costly (see further motivations in [54, Sec. 1.3]). Under this model linear operations with matrix-by-vector products can be performed with vectors.
Hereafter $a\ll b$ means that $a=o(b)$ but also that $b$ greatly exceeds $a$ in non-asymptotic context, $A^{+}$ denotes the Moore-Penrose pseudo inverse of $A$, and $m\geq n$; otherwise we can work with the $n\times m$ transpose $M^{T}$.
[…]
Computation of $M_{r}$ and $r$-top SVD involves order of $mn^{2}$ floating point operations (flops) and the memory space for about $mn$ scalars (see [15, Fig. 8.6.1]), although the output $r$-top SVD only uses memory for $(m+n+r)r$ scalars.
1.1.3. Variations and computational equivalence of LRA tasks. Next we state LRA task by generalizing this and other known representations of LRA as follows: given a matrix $M\in\mathbb{R}^{m\times n}$ and a pair of integers $r_{-}$ and $r_{+}$ such that
$0<r_{-}\leq r_{+}\ll n\leq m,$ (1.1)
fix two integers $k$ and $\ell$ such that
$k\ll m,~{}\ell\ll n,~{}r\leq\min\{k,\ell\},~{}{\rm and}~{}r_{-}\leq r\leq r_{+}$ (1.2)
and seek a rank-$r$ approximation of $M$ by a triplet of matrices $X\in\mathbb{R}^{m\times k}$, $Y\in\mathbb{R}^{k\times\ell}$, and $Z\in\mathbb{R}^{\ell\times n}$ such that the error matrix $M-XYZ$ has a small spectral and/or Frobenius norm. [^3] For $r_{-}=r_{+}$, relationships (1.2) define a target rank $r=r_{-}=r_{+}$. Otherwise, more realistically, they define a target range for rank $r$.
If the matrices $X=C$ and $Z=R$ are made up of the columns and rows of $M$, respectively, then $Y$ is commonly denoted by $U$ (or by $G$ by the followers of the pioneering work [17]) and the triplet $\{X,Y,Z\}$ is called CUR LRA (or CGR LRA) and is a particularly memory efficient LRA.
For any $r\leq\min\{k,\ell\}$ we can transform any LRA triplet $\{X,Y,Z\}$ for $M$ into its $r$-top SVD and further into its CUR (CGR) LRA by using $O(mk^{2}+n\ell^{2})$ flops, $(m+k)k+(n+\ell)\ell$ scalars and $O(mr^{2})$ flops, $(m+n+r)r$ scalars, respectively (see Appendices B and C.2, respectively).
1.1.4. LRA via random sampling. Impressive progress in LRA in the last decades relied on random sampling algorithms (see [8, 6, 26, 20, 56], and the references therein).
We need an efficient LRA algorithm for incorporation into our Alg. 2.1 of Sec. 2 and for natural candidates consider random sampling algorithms, such as the following algorithm of [54], which modifies those of [6, Thm. 4.7] and [6, Thm. 4.3, display 1].
Algorithm 1.1.
[LRA by means of random sampling].
Input:
An $m\times n$ black box matrix $M$, given implicitly by a subroutine for its multiplication by vectors, and positive integers $r_{-}$ and $r_{+}$ satisfying (1.1).
Output:
An integer $r$ satisfying (1.2) and, moreover, equal to $\min\{k,\ell\}$ and a triplet of matrices $\{X,Y,Z\}$ such that $XYZ$ is a rank-$r$ matrix close to $M$.
Initialization:
Fix a pair of integers $k$ and $\ell$ satisfying (1.2) and generate a pair of independent random test matrices $F\in\mathbb{R}^{\ell\times m}$ and $H\in\mathbb{R}^{n\times k}$.
Computations:
  1. 1.
    Compute the matrices $Z:=FM\in\mathbb{R}^{\ell\times n}$ and $MH\in\mathbb{R}^{m\times k}$, which are two sketches (samples) of $M$.
  2. 2.
    Compute the matrix $X\in\mathbb{R}^{m\times k}$ denoting the Q factor of the thin QR factorization of $MH$ – (cf. [15, Thm. 5.3.2]).
  3. 3.
    Compute the matrx $Y:=(FQ)^{+}\in\mathbb{R}^{k\times\ell}$.
  4. 4.
    Output the triplet of matrices $\{X,Y,Z\}$.
The algorithm involves $O((k+\ell)mn)$ flops [^4] One can do with just $O(mn\log(k+\ell))$ flops to ensure a little weaker bounds on failure probability [20, Sec. 11.2]. and about $(m+n)(k+\ell)$ scalars, not counting the cost of the generation of random matrices $F$ and $H$. Like various other random sampling LRA algorithms, Alg. 1.1 is universal, that is, devised for all matrices $M\in\mathbb{R}^{m\times n}$; for proper pairs of dense test matrices $f$ and $H$ and integers $k$ and $\ell$ in $O(r\log(r))$ it belongs to the class of relative error LRA algorithms: for any fixed pair of positive $\epsilon$ and $\delta$ it outputs rank-$r$ approximations of $M$ whose Frobenius error norm is within a factor of $1+\epsilon$ from the optimal value $||M-M_{r}||_{F}$ with a failure probability at most $\delta$ (see the bibliography cited above).
1.1.5. Superfast (sublinear cost) LRA. The LRA algorithms of [28, 4, 21, 23, 5] (randomized) and [24, Part III] (deterministic) are known to run superfast, aka at sublinear cost, that is, to use $o(mn)$ scalars and flops for a fixed pair of $\epsilon$ and $\delta$ [^5] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [34, 36], and we extend this concept to other matrices.
Unfortunately, already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [40, 41, 42, 43]). If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0. Likewise, any deterministic or randomized superfast estimation of the error norm of an LRA fails on the matrices of that small family. Hence a superfast algorithm cannot be universal for computation of LRA or even for its a posteriori error norm estimation.
Accordingly, the randomized LRA algorithms known to be superfast are restricted to some large and important classes of input matrices such as Symmetric Positive Semidefinite (SPSD) matrices, Distance matrices, or a subclass of Cauchy matrices. Furthermore, strictly speaking they are NOT superfast because they involve a posteriori error norm estimation in linear or superlinear time. The latter deficiency is only avoided in the deterministic algorithm of [24, Part III] for LRA of SPSD matrices – its output error norm is estimated as by-product, at no additional cost.
1.1.6. Conflicts #$\#1$ for LRA. Trying to optimize simultaneously various permissible parameters of LRA such as arithmetic cost and the costs of randomization, storage, and communication, representing the computational cost one runs into the conflict with the quality of a priori error bounds one (see [54, Sec. 3.9]), which boils down to the conflict between the two main goals of devising efficient LRA algorithms, that is, minimization of (a) the output error norm and (b) represented by the memory space and number of flops involved. Here are some specific samples of this omnipresent LRA conflict #$\#1$, routinely resolved by means of trade-off.
(i) The known random sampling LRA algorithms need unbounded memory size and flop number to ensure with a high probability (whp) the relative error norm bound within $1+\epsilon$ from optimal under the Frobenius norm for $\epsilon\rightarrow 0$ [20, Sec. 10.3]. Furthermore, the LRA algorithms supporting the same output error bound under the spectral rather than Frobenius norm are more costly (see [56, Ch. 6.2], [13, Thm. 3.4]).
(ii) Adaptive Cross-Approximation (ACA) iterations, [^6] We briefly recall their simplified version of Cross-Approximation (C-A) iterations in Appendix E. which are empirically superfast for a large class of matrices $M$, output LPAs whose error norm bound is at best within a factor of $2\sqrt{mn}$ from optimal under both spectral and Frobenius matrix norms (see [30, Cor. 2]).
(iii) Likewise, the output error norms of the relative error LRAs of [24, Part III] for $n\times n$ Symmetric Positive Semidefinite (SPSD) matrices $M$ may exceed the optimal spectral and Frobenius norms $||M-M_{r}||$ by a factor of $n$.
1.2 Our progress (briefly)
For matrices with fast decaying spectra of singular values (hereafter FDS matrices) our algorithms decrease the spectral norm of the output error matrix of the known LRA algorithms at sublinear additional cost, that is, superfast. Our numerical tests with both synthetic and real world matrices are in good accordance with these formal estimates. Furthermore, some variants of our entire algorithms run superfast, and we empirically observe that they still output accurate or even near-optimal LRAs for a large subclass of FDS matrices.
1.3 LRA by means of escalation
Next we outline our rather simple techniques.
1.3.1. More definitions and two theorems. Hereafter $\sigma_{j}(M)$ denotes the $j$th top (largest) singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$. $||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Sec. 5.
In particular we only recall the Eckart-Young-Mirsky theorem (cf. [50, 12, 27], [15, Thm. 2.4.8]) under the spectral norm:
Theorem 1.1.
$\sigma_{r+1}(M)=||M_{r}-M||=\min_{X:~{}\operatorname{rank}(X)\leq r}||M-X||$.
Theorem 1.2.
[15, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
$|\sigma_{j}(M+E)-\sigma_{j}(M)|\leq||E||~{}{\rm for}~{}j=1,\dots,\min\{m,n\}.$
1.3.2. Gaps in the spectrum and escalation of LRAs. For a fixed $r$ in a target range, we fix an upper rank $\rho$ satisfying $r\leq\rho\ll n$ and then (i) compute a rank-$\rho$ approximation of $M$ given by a triplet $\{X,Y,Z\}$ such that
$||M-M(\rho)||=\beta\sigma_{\rho+1}(M),~{}{\rm for}~{}M(\rho)=XYZ,~{}1\leq\beta\leq\beta_{+}~{}{\rm and~{}a~{}fixed~{}bound}~{}\beta_{+},$ (1.3)
and (ii) compute and output the $r$-top SVD of $M(\rho)$.
The relative error random sampling algorithms of [20, 54] use from $O(mn\log(\rho))$ to $O(mn\rho)$ flops and about $(m+n)\rho$ scalars to perform stage (i), while their heuristic variations in our Sec. 4 perform stage (i) at a lower cost; deterministic Alg. B.1 of Appendix B uses $O(m\rho^{2})$ flops and about $(m+n)\rho$ scalars to perform stage (ii). [^7] By applying the algorithm of [14] instead of Alg. B.1 at stage (ii), one can a little accelerate the computations while only slightly increasing the error norm bound.
In Alg. 2.1 of Sec. 2 we elaborate upon this outline and deduce that
$||M-(M(\rho))_{r}||\leq\sigma_{r+1}(M)+2\beta\sigma_{\rho+1}(M)=\Big{(}1+\frac{2\beta\sigma_{\rho+1}(M)}{\sigma_{r+1}(M)}\Big{)}\sigma_{r+1}(M).$ (1.4)
The output error norm $||M-M(\rho)_{r}||$ of Alg. 2.1 is small if the value $\frac{\sigma_{r+1}(M)}{2\beta\sigma_{\rho+1}(M)}$ is large, that is, if there is a large gap in the segment $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of the spectrum of $M$.
For comparison, a rank-$r$ approximation of $M$ is only interesting if the value $\frac{\sigma_{1}(M)}{\sigma_{r+1}(M)}$ is large, that is, if there is a gap in the leading segment $[\sigma_{1}(M),\sigma_{r+1}(M)]$ of the spectrum.
We can reveal gaps in the spectrum by approximating the values of $||M-(M(\rho))_{r}||$ for various pairs of $\rho$ and $r$ and comparing the output error norms.
1.3.3. Superfast variations. We can compute the $r$-truncation of $M(\rho)$ superfast for any matrix $M$ if $\rho$ is small such that $\rho^{2}=o(n)$, while some superfast variations of the known LRA algorithms (e.g., those of [20, 54] discussed in our in Sec. 4) output reasonably accurate rank-$\rho$ approximations $M(\rho)$ for a large class of FDS matrices $M$.
We cannot estimate the spectral error norm of a matrix superfast, and even its LAPACK heuristic estimation is not superfast, but we propose a heuristic superfast 1-norm estimator in Sec. 5 and test it in Sec. 6.4.
1.3.4. Iterative LRAs. In Alg. 3.1 of Sec. 3, which is a competitive alternative to Alg. 2.1, we write $X_{0}:=0$, $E_{i}:=M-X_{i}$, and $X_{i}:=(X_{i-1}+E_{i}(\rho))_{r}$ for $i=1,2,\dots,k$. We compress the matrices $X_{i-1}+E_{i}(\rho)$, that is, keep their ranks bounded by $r$ – by working with their $r$-truncation, and we do this superfast as long as their rank bound $\rho+r$ is low. For computation of the matrix $E_{i}(\rho)$ we use the algorithms of stage (i) of Alg. 2.1. This way we extend to LRA the popular iterative refinement algorithms.
The algorithm is essentially reduced to $k$ applications of Alg. 2.1.
In our tests Alg. 3.1 output LRAs with relative output errors about 1.1 for a matrix class marked Shaw for which Alg. 2.1 failed, but for all other inputs Alg. 3.1 was less accurate than Alg. 2.1 applied with the upper rank $\rho$ twice as large versus Alg. 2.1.
Our upper estimates for the accuracy of both algorithms are close to one anaother but are both pessimistic, and it seems that further comparative study of the two algorithms is in order.
For completeness, we cover iterative refinement of LRA with no compression in Appendix F, but unlike Alg. 3.1, we cannot provide any support for its numerical stability.
1.4 Related works
Motivated by exposition of the progress in LRA in [20, 54], we extend the rudimentary study of LRA escalation in [40, 41, 43, 42] (cf. the end of Sec. 4).
The algorithms in [25, 39] extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [51, Secs. 3.3.4 and 4.2.5], [19, Ch. 12 and Sec. 20.5], [7], [15, Secs. 3.5.3 and 5.3.8], [3, Secs. 1.4.6, 2.1.4, and 2.3.7]) but also widely applied in various other linear and nonlinear matrix computations (see [51, page 223 and the references on page 225]). Its extension to LRA is not straightforward – because the rank of the computed approximations grows recursively; we block this growth by applying re-compression, tracing such techniques to [31, 32, 37, 46, 38, 35] and [34, Secs. 6.3-6.5].
Very different iterative refinement of LRA in [25] uses no crude auxiliary LRA but instead bounds the Principal Angle Distances between singular spaces of $M$ and its LRAs and incorporates randomization based on the leverage scores of [8].
Historically, we first succeeded with iterative refinement of LRA and then noticed that escalation alone is quite competitive.
In our choice of abridged SRHT matrices for sparse multipliers we follow [40, 41, 42, 43], which proved some crude but reasonable probabilistic upper bounds on output errors of some superfast LRA algorithms under randomization in the input matrix space defined by a fixed probability structure in that space. [^8] These papers continued the study in [45, 44, 47], which introduced a probability structure in the space of input matrices to motivate application of Gaussian elimination without pivoting.
The relevance of the gaps in the spectrum of the singular values of a matrix to its LRA is well-known but has not been exploited in the way that we propose, test, and analyze.
Likewise, from the comments of an expert on an earlier draft of this paper, we learned that [9] also worked with error matrices of LRA, but in the directions quite different from ours.
1.5 Organization of our paper
We present our LRA escalators in Secs. 2 and 3 and discuss superfast computation of a crude LRA in Sec. 4. We specify a superfast heuristic 1-norm estimator in Sec. 5, cover our tests in Sec. 6, and devote our brief Sec. 7 to conclusions. In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails. In Appendix B we compute $r$-top SVD of an LRA or any low-rank matrix superfast. In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA. In Appendix D we generate abridged SRHT matrices. In Appendix E we briefly recall ACA (C-A) iterations. In Appendix F we iteratively refine LRA with no compression.
2 LRA by means of escalation
Next we elaborate upon escalation outlined in Sec. 1.3.2.
Algorithm 2.1.
LRA by means of escalation.
Input:
Two positive integers $m$ and $n\leq m$, an $m\times n$ matrix $M$, possibly given implicitly by its sketches, three positive values $\widehat{\tau}$, $\widehat{\mu}$, and $r$, $r\ll n$, denoting bound on the overall computational cost, [^9] Here and hereafter we use a single character $\tau$ to represent both flops and memory size involved. an absolute error norm bound, and a fixed integer in a target range, respectively; three auxiliary Algs. $\mathbb{A}$, $\mathbb{B}$, and $\mathbb{C}$ which for any $\rho$, $r\leq\rho\leq n$, compute a (crude but reasonably close) rank-$\rho$ approximation of $M$, compute $r$-top SVD of a rank-$\rho$ matrix of size $m\times n$, and estimate the norm of the difference of $M$ and another $m\times n$ matrix, respectively, by performing within computational cost bounds $\tau(\rho)$, $\tau(\rho,r)$, and $\tau^{\prime}$, respectively. [We elaborate upon Algs. ${\bf\mathbb{A}}$, ${\bf\mathbb{C}}$, and ${\bf\mathbb{B}}$ in Secs. 4 and 5 and Appendix B, respectively.]
Output:
FAILURE or SVD of a rank-$r$ matrix $X\approx M$ and the norm $||M-X||$ not exceeding $\widehat{\mu}$.
Initialization:
Fix the maximal upper rank $\rho$ satisfying the following bounds (see Remark 2.1),
$\tau(\rho)+\tau(\rho,r)+\tau^{\prime}\leq\widehat{\tau}~{}{\rm and}~{}r\leq\rho\ll n\leq m.$ (2.1)
Computations:
  1. 1.
    Apply Alg. $\mathbb{A}$ to output a rank-$\rho$ approximation of $M$ given by SVD of some matrix $M(\rho)$.
  2. 2.
    Apply Alg. ${\bf\mathbb{B}}$ to that matrix to compute its $r$-top SVD,101010 Alg. B.1 of Appendix B performs this stage superfast where $\rho^{2}\ll n$. which is SVD of $X:=M(\rho)_{r}$.
  3. 3.
    If Alg. ${\bf\mathbb{C}}$ shows that
    $\mu:=||E||\leq\widehat{\mu}~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
    then output that SVD and the estimated norm $\mu$. Otherwise output FAILURE.
Remark 2.1.
[The computational cost bound, the choice of an upper rank, and a simplified version of Alg. 2.1.] (i) Bound (2.1) implies that the overall computational cost $\tau(\rho)+\tau(\rho,r)+\tau^{\prime}$ stays within tolerance $\widehat{\tau}$. (ii) Only limited by Eqn. (2.1), we maximize the upper rank $\rho$ – to decrease the error norm bound $\mu$. (iii) By increasing the upper bound $\widehat{\mu}$ on the output error norm $\mu$, we can decrease the cost of computations at stage 2, which, however, tends to be dominated. (iv) In Sec. 6.3 we test a simplified version of Alg. 2.1 where we impose no upper bounds $\widehat{\tau}$ and $\widehat{\mu}$ and display the output error norms for $\rho:=ir$ and $i=2,3,4,5$.
Next we prove bound (1.4).
Theorem 2.1.
For $\mu$ of Eqn. (2.2) and $\beta$ of (1.3), it holds that
$\mu\leq\sigma_{r+1}(M)+\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||=\beta\sigma_{\rho+1}(M).$ (2.3)
Proof.
Represent $E$ of (2.2) as $M-M_{r}+M_{r}-M(\rho)_{r}$, apply triangle inequality, and obtain
$||E||\leq||M-M_{r}||+||M_{r}-M(\rho)_{r}||=\sigma_{r+1}(M)+||M_{r}-M(\rho)_{r}||.$
Thm. 1.2 implies that $||M_{r}-M(\rho))_{r}||\leq\nu$. Combine the latter two bounds. ∎
Corollary 2.1.
[Relative output error norms of Algs. $\mathbb{A}$ and 2.1.] Write (cf. (1.3))
$\alpha:=\alpha(M,\rho,r):=\frac{||E||}{\sigma_{r+1}(M)}-1~{}{\rm and}~{}\beta:=\beta(M,\rho):=\frac{\nu}{\sigma_{\rho+1}(M)}.$ (2.4)
Then the rank-$r$ approximation of $M$ output by Alg. 2.1, for a fixed pair of $r$ and $\rho$, is optimal up to within the absolute and relative error norms $\nu$ and $1+\alpha$, respectively, such that
$1+\alpha\leq\frac{\nu}{\sigma_{r+1}(M)}=\frac{\beta\sigma_{\rho+1}(M)}{\sigma_{r+1}(M)}.$ (2.5)
Remark 2.2.
How sharp are the upper bounds (2.3)-(2.5), that is, how much can we lose when we apply triangle inequality and Thm. 1.2 in the proof of Thm. 2.1? The loss can be significant: substitute $\rho=r$ into (2.3) and obtain $\mu:=||M-M(\rho)_{r}||\leq(1+\beta)\sigma_{r+1}(M)$, but more directly obtain $\mu:=||M-M(\rho)_{r}||=||M-M(r)||=\beta\sigma_{r+1}(M)$ because $M(\rho)_{r}=M(r)$.
3 Iterative escalation with compression
Next we elaborate upon iterative escalation with compression, sketched in Sec. 1.3.4.
Algorithm 3.1.
Iterative escalation.
Input and OUTPUT
as in Alg. 2.1, except that the input of Alg. 3.1 additionally includes a positive integer $k$ and an Alg. $\mathbb{S}$ that adds or subtracts a pair of black box $m\times n$ matrices at a computational cost $\tau(\pm)$.
Initialization:
Write $i:=0$, $X_{0}:=0$, and $E_{0}:=M$ and fix an upper rank $\rho$ satisfying
$2r\leq\rho\leq n\leq m,~{}~{}(\tau(\rho)+\tau(\rho,r)+\tau(r^{\prime})+2\tau(\pm))~{}k\leq\widehat{\tau}.$ (3.1)
Computations:
  1. 1.
    Apply Alg. ${\bf\mathbb{A}}$ to the matrix $E_{i}$ and the upper rank $\rho$. [The algorithm outputs a rank-$\rho$ approximation of $E_{i}$ given by SVD of a matrix $E_{i}(\rho)$.]
  2. 2.
    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$.
  3. 3.
    Apply Alg. ${\bf\mathbb{B}}$ to this matrix and the fixed $r$ to compute $r$-top SVD of $M_{i}(\rho)$, that is, SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$.
  4. 4.
    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if Alg. ${\bf\mathbb{C}}$ shows that $2\mu_{i}:=2\mu_{i}(M,\rho)\leq\widehat{\mu}$.
    Otherwise output FAILURE if $i=k-1$.
  5. 5.
    Otherwise compute $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1.
Remark 3.1.
[The computational cost bound, the choice of an upper rank, and a simplified version of Alg. 3.1.] (i) (3.1) implies that the overall computational cost $\tau(\rho)+\tau(\rho,r)+\tau^{\prime}$ stays within tolerance $\widehat{\tau}$. (ii) As in Alg. 2.1 we maximize the upper rank $\rho$ to decrease the output error norm, being only limited by a bound on the overall computational cost. (iii) In principle, we can allow an upper rank $\rho$ to vary as $i$ varies; then we can readily extend the analysis of the algorithm. (iv) In Sec. 6.3 we test a simplified version of Alg. 3.1 where we impose no upper bounds $\widehat{\tau}$ and $\widehat{\mu}$, fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ elsewhere.
In the following extension of Thm. 2.1 we estimate the output error norm of Alg. 3.1.
Theorem 3.1.
For the error matrices $E_{i+1}$ of Alg. 3.1, $\nu_{i}:=||E_{i}-E_{i}(\rho)||$, $i=0,1,\dots,k-1$, and an upper bound
$\beta\geq\beta_{i}(M,\rho):=\frac{\nu_{i}}{\sigma_{\rho+1}(E_{i})}~{}{\rm for}~{}i=0,1,\dots,k-1$ (3.2)
on the relative errors of Alg. $\mathbb{A}$ (unlike (1.3) we write $\beta$ rather than $\beta_{+}$), it holds that
$\mu_{i+1}:=||E_{i+1}||\leq||M-M_{r}||+\nu_{i}=\sigma_{r+1}(M)+\beta\sigma_{\rho+1}(E_{i}).$
Proof.
Recall from Alg. 3.1 that $E_{i+1}=M-X_{i+1}$ for $X_{i+1}=(X_{i}+E_{i}(\rho))_{r}$. Hence $E_{i+1}=M-M_{r}+M_{r}-X_{i+1}$, and so $||E_{i+1}||\leq||M-M_{r}||+||M_{r}-X_{i+1}||=\sigma_{r+1}(M)+||M_{r}-X_{i+1}||$.
It remains to prove that
$||M_{r}-X_{i+1}||\leq\beta\sigma_{\rho+1}(E_{i}).$ (3.3)
Since $X_{i+1}=(X_{i}+E_{i}(\rho))_{r}$, it follows that $M_{r}-X_{i+1}=M_{r}-(X_{i}+E_{i}(\rho))_{r}$. Combine this equation with Thm. 1.2 and obtain that $||M_{r}-X_{i+1}||\leq||M-(X_{i}+E_{i}(\rho))||=||E_{i}-E_{i}(\rho)||$. Now bound (3.3) and hence the theorem follow from Eqn. (3.2). ∎
Corollary 3.1.
Extend Cor. 2.1 by writing $\alpha_{i+1}$ and $E_{i+1}$ instead of $\alpha$ and $E$. Then
$\alpha_{i+1}:=\alpha_{i+1}(M,r,\rho):=\frac{||E_{i+1}||}{\sigma_{r+1}(M)}-1\leq\beta\frac{\sigma_{\rho+1}(E_{i})}{\sigma_{r+1}(M)},~{}i=0,1,\dots,k-1.$ (3.4)
Lemma 3.1.
Recall that $\sigma_{j}(X)=0$ for $j>n$ and write $\sigma_{j}(X)=\infty$ for any $m\times n$ matrix $X$ and $m\geq n$. Then $\sigma_{j+r+1}(M)\leq\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$ for $E_{i}$ of Alg. 3.1 and all $i$ and $j$.
Proof.
Recall that $\operatorname{rank}(M-X_{i})\leq r$ for all $i\geq 0$, $\sigma_{j+1}(E_{i})$ is an optimal rank-$j$ approximation of the matrix $E_{i}$ by virtue of Thm. 1.1, and rank-$\rho$ approximation of $M$ with $Y$ does not exceed rank-$(\rho-r)$ approximation of $M$ with $Y+X$ if $\operatorname{rank}(X)=r$. Combine these observations. ∎
Combine bound (3.4) with the lemma for $j=\rho+1$ to obtain
Corollary 3.2.
The singular values $\sigma_{\rho+1}(E_{i})$ of Eqn. (3.4) satisfy
$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,k-1.$ (3.5)
Remark 3.2.
[Error norm bounds.] For $E_{i+1}=E$ and $E_{i}=M$, bound (3.4) turns into bound (2.5) of Cor. 2.1, and this is precisely what we have at the initial stage for $i=0$. For $i>0$ bound (3.4) can be weaker or stronger than that, but it cannot be stronger than (2.5) with $\rho$ replaced by $\rho+r$. In our tests for Algs. 2.1 and 3.1 in Sec. 6.3 for for matrices Shaw and $r=20$ Alg. 3.1 for $\rho=2r$ superseded Alg. 2.1 for $\rho=ir$, $i=3,4,5$. This implies that the upper bound of Cor. 3.2 is overly pessimistic for worst case inputs, similarly to bound of Cor. 2.1 (cf. Remark 2.2).
4 LRA by means of random sampling with sparse multipliers
Alg. 1.1 is near-optimal with a high probability (whp) but not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT [20], [54, Sec. 3.9].
One can accelerate it to a superfast level by using Ultra-Sparse Rademacher test matrices [54, Sec. 3.9]), [6], which depend on an integer parameter $s$, defining sparsity. They are sparse and support superfast LRAs for smaller values $s$.
Likewise, in Appendix D we turn SRHT test matrices into abridged SRHT test matrices of a small length, which support superfast LRA; this recipe can be extended to various other recursively generated test matrices such as SRFT matrices and those of [20, Remark 4.6], [15, Sec. 5.1.8].
According to [22], such acceleration tends to make the output LRAs somewhat less reliable. Apparently for this reason, our tests for synthetic inputs from [55, Sec. 7.3.1] failed with abridged SRHT test matrices of length 3, while the same tests succeeded with Gaussian test matrices. For the other inputs, however, our tests succeeded about equally well with test matrices of both kinds, even where the output LRAs of Alg. $\mathbb{A}$ were much cruder with abridged SRHT test matrices than with Gaussian random ones.
[40, 41, 42, 43] studied primitive algorithms for CUR LRAs (cf. Appendix C.1) generated by pairs of test matrices made up of two random subsets of $k$ rows and $\ell$ columns of the identity matrix. These algorithms run at very low cost and output rather crude LRAs, and so [40, 41, 42, 43] proposed to strengthen them with escalation, by computing the sets of $k$ rows and $\ell$ columns as proper subsets of the sets of $p>k$ random rows and $q>\ell$ random columns of the identity matrix. [40, 41, 42, 43] call these algorithms cynical – representing a minor advance from the primitive ones. Not competitive for LRA, such cynical algorithms were predecessors of our use of escalation and could still be considered for Alg. $\mathbb{A}$.
5 Superfast 1-norm estimator
Next we propose a superfast variation of the LAPACK estimator for 1-norm $||E||_{1}$ of a matrix $E\in\mathbb{R}^{m\times n}$ (cf. [19, Alg. 15.3], [51, Alg. 5.3.1]) and recall that $||E||_{1}=||E^{T}||_{\infty}$, $||E||_{\infty}\leq n||E||$, $||E||_{1}\leq m||E||$, and $||E||\leq||E||_{1}||E||_{\infty}$ (see [15, Eqns. (2.3.11), (2.3.12) and Cor. 2.3.2])).
Empirically, the LAPACKl norm-1 estimator converges in about 4.5 iterations on the average and is superfast at every iteration, except for its initial and verification stages, at which $E$ is multiplied by dense auxiliary vectors. We accelerate these stages by using sparse vectors instead. We keep a small number of randomly sampled components of the original vectors, set the remaining components to 0, and scale the resulting vectors to turn their 1-norm into 1; we only test our 1-norm estimator for matrices $E$ representing the errors of LRA.
Algorithm 5.1.
[A superfast 1-Norm Estimator.]
Input:
An $m\times n$ matrix $E$, a positive integer $k\ll n\leq m$, and a positive tolerance TOL.
Output:
A vector ${\bf u}$ such that $||E{\bf u}||_{1}\approx||E||_{1}$.
Initialization:
Generate a sparse unit vector ${\bf v}$ such that $||{\bf v}||_{1}=1$ as follows: choose $k$ random integers $i_{1},\dots,i_{k}$ in the range from 1 to $n$ and let ${\bf v}=\frac{1}{k}(v_{i})_{i=1}^{n}$ where $v_{i_{j}}=1$ for $j=1,\dots,k$ and $v_{i}=0$ for all other $i$. Write COUNTER:=0.
Computations:
  1. 1.
    Compute the vectors ${\bf u}=E{\bf v}$, ${\bf w}=(\operatorname{sign}(u_{i}))_{i=1}^{n}$, and ${\bf x}=E^{T}{\bf w}$.
  2. 2.
    If $||{\bf x}||_{\infty}\leq||{\bf u}||_{1}$, output ${\bf u}$ and terminate.
  3. 3.
    Choose $j$ such that $|x_{j}|=||{\bf x}||_{\infty}$ and let ${\bf v}={\bf e}_{j}$. Increase COUNTER by 1. Go to stage 1 if COUNTER$<$Tol. Otherwise stop and output FAILURE.
We also apply the algorithm to a small number of other choices of the initial sparse unit vector ${\bf v}$ and output ${\bf u}$ that maximizes the norm $||E{\bf u}||_{1}$. For these choices we first define vectors of dimension $n$ sharing $k$ components on $k$ random positions with the following vector
$\left(1,-1-\frac{1}{n-1},1+\frac{2}{n-1},1-\frac{3}{n-1},\dots,(-1)^{n-1}2\right)^{T}$
and otherwise filled with 0s; then scale them to make their 1-norm equal to 1.
6 Numerical experiments
We devote the next three subsections to the tests of our LRA algorithms on synthetic and real-world input matrices devote Sec. 6.4 to testing Alg. 5.1.
We implemented our algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory. We called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
6.1 Input Matrices
[…]
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [20, Sec. 7.1] for more details).
We display the distribution of the top 50 singular values of these three matrices in Fig. 1.
We set the target ranks $r$ equal to 20 for Shaw, 45 for Gravity, and 11 for SLP matrices.
Figure 1: Spectra of singular values of some real world matrices
Furthermore, we computed rank-$r$ approximations for the three matrices, StreamVel, MinTemp, and MaxCut, used in numerical experiments in [55, Sec. 7.3.2].
The $10738\times 5001$ matrix StreamVel is involved in numerical simulation (DNS) on a coarse mesh of the 2D Navier–Stokes equations for a low-Reynolds number flow around a cylinder.
The $19264\times 7305$ matrix MinTemp tabulates meteorological variables at weather stations across the northeastern United States on days during the years 1981–2016.
The $2000\times 2000$ matrix MaxCut provides a high-accuracy solution to the MaxCut SDP for a sparse graph [18].
In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2 and set the target rank $r$ equal to 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
Synthetic Input Matrices.
We generated random synthetic $1024\times 1024$ input matrices of five classes, two of them – with fast and slowly decaying spectra – as the products $U\Sigma V^{T}$, with the matrices $U$ and $V$ of the left and right singular vectors of a Gaussian random matrix, respectively.
By letting $\Sigma=\textrm{diag}(v)$, $v_{i}=1$ for $i=1,2,3,\dots,20$, $v_{i}=\frac{1}{2^{i-20}}$ for $i=21,\dots,100$, and $v_{i}=0$ for $i>100$, we arrived at the matrices with fast decaying spectra.
By letting $\Sigma=\textrm{diag}(u)$, $u_{i}=1$ for $i=1,2,\dots,20$, and $u_{i}=\frac{1}{(1+i-20)^{2}}$ for $i>20$, we arrived at the matrices with slowly decaying spectra.
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the next three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [55, Sec. 7.3.1], which involved effective rank parameter $R=20$.
(i) Low-rank + noise matrices
${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$
for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, also involves parameter $\xi$, equal to $10^{-4}$ for the Low Rank Low Noise matrix, to $10^{-2}$ for the Low Rank Med Noise matrix, and to $10^{-1}$ for the Low Rank High Noise matrix.
(ii) We generated Polynomial Decay matrices
${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p}),$
where $p=0.5$ for Poly Decay Slow, $p=1$ for Poly Decay Med, $p=2$ for Poly Decay Fast.
(iii) We generated Exponential Decay matrices
${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q}),$
with $q=0.01$ for Exp Decay Slow, $q=0.1$ for Exp Decay Med, $q=0.5$ for Exp Decay Fast.
6.2 Two options for Alg. $\mathbb{A}$
In our tests of Algs. 2.1 and 3.1 we used the following two options for Alg. $\mathbb{A}$, which computes a crude rank-$\rho$ approximation of a matrix $M$.
(i) Fast (although not superfast) randomized algorithm of [54] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times(\rho)$.
(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{R}^{(2\rho)\times 1024}$ and $H\in\mathbb{R}^{1024\times\rho}$ replaced Gaussian ones.
6.3 Test results for Algs. 2.1 and 3.1
In our tests of escalation Alg. 2.1 we output the relative error norms
$\text{err}=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$
[…]
In our tests of iterative escalation Alg. 3.1 we output the relative error norms
[…]
We only display the error norms output in the first three iterations of Alg. 3.1 and notice that, except for the matrices representing high noise or slow decay, the second iteration greatly decreased the output error norms, while the third iteration tended to keep the error norms at more or less the same level.
[…]
For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple of $r$, $\rho$, and an input matrix type for Alg. 2.1, and per the pair of $r$ and an input matrix type for Alg. 3.1, except that we ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger size.
Our tables display the mean of the relative error norms observed in these tests.
For Alg. 2.1 we also display the standard deviation.
In Tables 6.1, 6.2, 6.5, and 6.6 we display the results of our tests with both Gaussian random and abridged SRHT test matrices. This covers the cases where the output error norms were more or less in the same range for both test matrices (and usually quite similar in both cases, with only a slightly better results with Gaussian random test matrices).
Displaying the results of our tests with the synthetic input matrices from [55, Sec. 7.3.1] in Tables 6.3 and 6.4, however, we only show the test results with Gaussian random test matrices because all tests with abridged SRHT test matrices have failed, that is, have output with large errors. The tests with Gaussian random test matrices have output accurate LRAs for the same input matrices, except for those representing high noise or slow decay.
Alg. 2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT test matrices. This could have been predicted because the spectrum of the singular values $\sigma_{j}(M)$ of these matrices $M$ is flat for $j>20$ (decreasing roughly from $10^{-13}$ to between $10^{-15}$ and $10^{-16}$); in Table 6.1 we display the test results for the matrix Shaw and target rank $r=19$, which are about as favorable as for Gravity and Slow Decay.
In both cases of success and failure of both algorithms, our test results were in good accordance with our formal study. The relative error norms output by both Algs. 2.1 and 3.1 have consistently decreased and frequently to the optimal value 1 (up to within the measurement errors) as the spectra of the input matrices decayed. The error norms have also decreased as we increased the upper rank $\rho$ for Alg. 2.1 and, as we said already, in the transition from the first to the second iteration of Alg. 3.1.
The error norm bounds of Alg. 2.1 for $\rho=4r$ were consistently smaller than those of Alg. 3.1 in all three iterations, except for tests with the matrix Shaw for $r=20$. This exception implies that our estimate of Cor. 3.2 is overly pessimistic for worst case inputs, as we explained in Remark 3.2.
Input Matrix Multiplier $\rho$
$2r$ $3r$ $4r$ $5r$
Gravity Abr. SRHT 1.000 $\pm$ 1.036E-05 1.000 $\pm$ 7.815E-06 1.000 $\pm$ 7.888E-06 1.000 $\pm$ 6.683E-06
Gaussian 1.000 $\pm$ 7.495E-06 1.000 $\pm$ 5.211E-06 1.000 $\pm$ 7.214E-06 1.000 $\pm$ 5.432E-06
SLP Abr. SRHT 1.970 $\pm$ 4.083 1.000 $\pm$ 0.002 1.000 $\pm$ 1.018E-07 1.000 $\pm$ 1.362E-10
Gaussian 1.001 $\pm$ 0.0008160 1.000 $\pm$ 1.528E-07 1.000 $\pm$ 1.075E-10 1.000 $\pm$ 6.350E-14
Shaw, $r=19$ Abr. SRHT 1.000 $\pm$ 1.778E-06 1.000 $\pm$ 1.359E-06 1.000 $\pm$ 1.665E-06 1.000 $\pm$ 1.441E-06
Gaussian 1.000 $\pm$ 2.256E-06 1.000 $\pm$ 1.685E-06 1.000 $\pm$ 1.260E-06 1.000 $\pm$ 1.427E-06
Fast Decay Abr. SRHT 1.000 $\pm$ 3.086E-12 1.000 $\pm$ 3.590E-16 1.000 $\pm$ 2.470E-16 1.000 $\pm$ 2.617E-16
Gaussian 1.000 $\pm$ 6.651E-12 1.000 $\pm$ 5.117E-16 1.000 $\pm$ 4.552E-16 1.000 $\pm$ 4.590E-16
Slow Decay Abr. SRHT 1.000 $\pm$ 3.604E-05 1.000 $\pm$ 1.244E-06 1.000 $\pm$ 1.543E-07 1.000 $\pm$ 3.754E-08
Gaussian 1.000 $\pm$ 4.130E-05 1.000 $\pm$ 1.380E-06 1.000 $\pm$ 1.783E-07 1.000 $\pm$ 3.708E-08
Table 6.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, synthetic matrices with fast decaying spectra (Fast Decay), and synthetic matrices with slow decaying spectra (Slow Decay).
Input Matrix Multiplier 1st Itr. 2nd Itr. 3rd Itr.
Fast Decay Abr. SRHT 3.155000 1.0000 1.0000
Gaussian 3.1202 1.0000 1.0000
Slow Decay Abr. SRHT 5.0468 1.0003 1.0001
Gaussian 5.0755 1.0002 1.0001
Shaw for $r=20$ Abr. SRHT 2.8820E+01 1.0983 1.1225
Gaussian 1.8235E+01 1.1517 1.1189
Gravity Abr. SRHT 1.5762E+01 1.0000 1.0000
Gaussian 1.2917E+01 1.0000 1.0000
SLP Abr. SRHT 1.0931E+02 1.0014 1.0000
Gaussian 5.2205 1.0000 1.0000
Table 6.2: Test results for Alg. 3.1 for $\rho=2r$ on synthetic matrices with fast decaying spectra (Fast Decay), synthetic matrices with slow decaying spectra (Slow Decay), Shaw, Gravity, and discretized single layer potential operator (SLP).
Input Matrix $\rho$
$2r$ $3r$ $4r$ $5r$
Low Rank Low Noise 1.0416 $\pm$ 8.9977E-02 1.$0$$\pm$ 2.1197E-06 1.$0$$\pm$ 2.4039E-06 1.$0$$\pm$ 2.2834E-06
Low Rank Med Noise 1.4335 $\pm$ 1.7048E-01 1.0382 $\pm$ 3.1809E-02 1.0057 $\pm$ 1.4442E-03 1.0026 $\pm$ 5.8650E-04
Low Rank High Noise 5.6972 $\pm$ 8.6182E-01 4.8401 $\pm$ 4.3819E-01 4.0328 $\pm$ 2.3493E-01 3.7893 $\pm$ 2.1626E-01
Poly Decay Slow 2.0588 $\pm$ 1.8783E-01 1.6525 $\pm$ 2.0027E-01 1.3617 $\pm$ 8.6188E-02 1.2062 $\pm$ 8.6352E-02
Poly Decay Med 1.5384 $\pm$ 2.1907E-01 1.0315 $\pm$ 2.6585E-02 1.0028 $\pm$ 1.0190E-03 1.0009 $\pm$ 4.5652E-04
Poly Decay Fast 1.3133 $\pm$ 1.5431E-01 1.0001 $\pm$ 1.3887E-04 1.$0$$\pm$ 3.7489E-06 1.$0$$\pm$ 2.4495E-07
Exp Decay Slow 2.8587 $\pm$ 3.3389E-01 2.2772 $\pm$ 2.0481E-01 1.8244 $\pm$ 1.0970E-01 1.5721 $\pm$ 1.0528E-01
Exp Decay Med 1.5576 $\pm$ 1.2324E-01 1.0414 $\pm$ 4.9018E-02 1.0001 $\pm$ 9.0188E-05 1.$0$$\pm$ 3.7953E-07
Exp Decay Fast 1.3121 $\pm$ 1.4989E-01 1.$0$$\pm$ 6.4663E-11 1.$0$$\pm$ 3.6020E-16 1.$0$$\pm$ 3.0986E-16
Table 6.3: Test results for Alg. 2.1 on the synthetic input matrices from [55, Sec. 7.3.1] using Gaussian multipliers.
Input Matrix 1st Itr. 2nd Itr. 3rd Itr.
Low Rank Low Noise 1.3940 1.0000 1.0000
Low Rank Med Noise 1.4752 1.0386 1.0375
Low Rank High Noise 1.4507 1.5135 1.5332
Poly Decay Slow 1.5920 1.4154 1.4073
Poly Decay Med 1.5569 1.0345 1.0306
Poly Decay Fast 1.3784 1.0001 1.0002
Exp Decay Slow 1.5202 1.4956 1.4750
Exp Decay Med 1.4946 1.0115 1.0164
Exp Decay Fast 1.4235 1.0000 1.0000
Table 6.4: Test results for Alg. 3.1 on the synthetic input matrices from [55, Sec. 7.3.1] using Gaussian multipliers.
Input Multiplier $\rho$
Matrix Type $2r$ $3r$ $4r$ $5r$
StreamVel Abr. SRHT 1.5840 $\pm$ 1.2583 1.4736 $\pm$ 1.4538 1.0022 $\pm$ 2.6478E-03 1.0002 $\pm$ 2.7733E-04
Gaussian 1.0932 $\pm$ 5.8887E-02 1.0064 $\pm$ 2.9809E-03 1.0013 $\pm$ 1.4141E-03 1.0003 $\pm$ 3.1762E-04
MinTemp Abr. SRHT 2.4575 $\pm$ 2.5748E-01 1.6980 $\pm$ 2.3998E-01 1.2919 $\pm$ 1.3606E-01 1.1181 $\pm$ 9.7415E-02
Gaussian 2.4393 $\pm$ 3.0823E-01 1.7216 $\pm$ 2.2510E-01 1.3793 $\pm$ 7.6629E-02 1.1238 $\pm$ 8.3331E-02
MaxCut Abr. SRHT 1.0024 $\pm$ 1.3630E-03 1.0005 $\pm$ 1.6875E-04 1.0002 $\pm$ 6.8092E-05 1.0001 $\pm$ 2.9040E-05
Gaussian 1.0026 $\pm$ 1.3856E-03 1.0005 $\pm$ 2.1129E-04 1.0002 $\pm$ 6.6260E-05 1.0001 $\pm$ 2.4207E-05
Table 6.5: Test results for Alg. 2.1 on the real world input matrices from [55, Sec. 7.3.2]
Input Matrix Multiplier Type 1st Itr. 2nd Itr. 3rd Itr.
StreamVel Abr. SRHT 2.9700 1.0739 1.0300
Gaussian 2.8724 1.0397 1.0177
MinTemp Abr. SRHT 2.3340 1.5961 1.5079
Gaussian 2.3120 1.6410 1.4950
MaxCut Abr. SRHT 3.6200E+01 1.0312 1.0182
Gaussian 2.6202E+01 1.0302 1.0176
Table 6.6: Test results for Alg. 3.1 for $\rho=2r$ on real world input matrices from [55, Sec. 7.3.2].
6.4 Test results for Alg. 5.1
For each $n\times n$ matrix $M$, we computed its rank-10 approximation $\tilde{M}$ by its truncated SVD and then obtained the error matrix $E=M-\tilde{M}$. We generated initial vectors ${\bf v}$ according to the following
INITIALIZATION RULE: Define a vector ${\bf v}$ of length $n$ with coordinates $1/k$ at $k$ positions sampled at random independently under the uniform probability distribution; fill the vector with 0s elsewhere; do this for $k=1,\log\log(n),\log(n)\text{ and }n$.
We performed 100 tests for each pair of input matrix type and integer $k$ and measured the accuracy of the estimates by the ratio
$\text{accuracy}=\frac{||E||_{1}}{\text{est}},$
where est denoted the estimate of $||E||_{1}$ returned by the algorithm.
The ratio took on its minimal value 1 where the estimate est was optimal.
In our test most of the norm estimates est returned by the algorithm were within a factor of 2 from the optimal norm $||E||_{1}$, even where the initial vector had just a single nonzero entry.
All our tests converged within 6 iterations, regardless of the density of the initial vector.
7 Conclusions
We proposed, tested, and analyzed some simple but novel LRA algorithms using escalation and proved that at a dominated or even sublinear additional computational cost they significantly decrease the spectral norms of the output error matrices of the known algorithms applied to FDS matrices.
The results of our numerical tests with a variety of synthetic and real world matrices are in good accordance with our formal estimates.
Furthermore, some superfast (sublinear cost) variants of the known algorithms are known to output LRA that tend to be less accurate. In combination with our algorithms, however, this deficiency has been fixed in most of our tests, and it is a challenge to test the output accuracy of these and other superfast (sublinear cost) LRA algorithms for a greater variety of inputs.
Towards enhancing our progress to a larger class of matrices $M$, we nontrivially extended to LRA the popular technique of iterative refinement. So far the resulting algorithm was superior to our escalation LRA for only a single input class of matrices among 17 classes tested, but further comparative study may still be in order.
Another natural but yet unexplored challenge is to compare and possibly combine the iterative refinement LRA algorithm of [25] with our present escalation and iterative refinement algorithms.
Appendix
Appendix A Small families of matrices that are hard for superfast LRA
Any superfast algorithm fails to compute a close LRA of the following small families of matrices.
Example A.1.
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Furthermore, the same argument shows that superfast estimation of the norm of a matrix fails already on the same matrix family and that for both problems of LRA and norm estimation randomized superfast algorithms fail with error probability not close to 0.
Remark A.1.
How representative are the above matrix families of hard inputs? The matrices of these families represent data singularities and are not relevant to a large class of input matrices representing regular processes or, say, smooth surfaces. Furthermore, Example A.1 does not apply to various important special classes of input matrices such as symmetric or diagonally dominant.
Appendix B Computation of $r$-top SVD of a matrix product
For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [31, 32, 37, 46, 38]) based on computation of the SVDs of $A$ and $B$.
Algorithm B.1.
[Computation of an $r$-top SVD of a matrix product.]
Input:
Four integers $r$, $k$, $m$, and $n$ such that $0<r\leq k\leq\min\{m,n\}$ and two matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$.
Output:
Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{R}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
Computations:
  1. 1.
    Compute SVDs $A=U_{A}\Sigma_{A}V^{*}_{A}$ and $B=U_{B}\Sigma_{B}V^{*}_{B}$ where $U_{B}\in\mathbb{C}^{m\times k}$, $V_{B}^{*}\in\mathbb{C}^{k\times n}$, and $\Sigma_{A},V^{*}_{A},U_{B},\Sigma_{B}\in\mathbb{C}^{k\times k}$.
  2. 2.
    Compute $k\times k$ matrices $W=\Sigma_{A}V^{*}_{A}U_{B}\Sigma_{B}$, $U_{W}$, $\Sigma_{W}$, and $V^{*}_{W}$ such that $W=U_{W}\Sigma_{W}V_{W}^{*}$ is SVD, $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$, and $\Sigma=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ and $\Sigma^{\prime}=\operatorname{diag}(\sigma_{j})_{j=r+1}^{k}$ are the matrices of the $r$ top (largest) and the $k-r$ trailing (smallest) singular values of the matrix $W$, respectively. Output the matrix $\Sigma$.
  3. 3.
    Compute and output the matrices $U$ and $V$ made up of the first $r$ columns of the matrices $U_{A}U_{W}$ and $V_{B}V_{W}$, respectively.
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [15, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
Appendix C CUR decomposition, CUR LRA, and its link to top SVD
C.1 CUR decomposition and CUR approximation
Next we briefly revisit CUR LRA, which is particularly memory efficient. For five integers $k$, $\ell$, $m$, $n$, and $r$ satisfying (1.2), an $m\times n$ matrix $M$, and two sets $\mathcal{I}\subseteq\{1,\dots,m\}$ and $\mathcal{J}\subseteq\{1,\dots,n\}$ of cardinalities $k$ and $\ell$, respectively, define the submatrices
$R:=M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},C:=M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{}{\rm and}~{}G:=M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}$
of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$ (cf. [8]), although we recall notation $G$ for nucleus and CGR for the factorization in [30], originated with the decomposition itself in the 1990s [17]. In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
Given a $k\times\ell$ generator submatrix $G$ of $M$, we can compute CUR approximation of $M$ by using $k\ell$ scalars and $O(k\ell\min\{k,\ell\})$ flops.
Theorem C.1.
(See [43].) For matrices $C$, $U$, and $R$ defined above, $M=CUR$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large, but empirically the Cross-Approximation (C-A) celebrated iterations, briefly recalled in Appendix E, tend to converge fast to quite an accurate rank-$r$ CUR approximation for a large class of matrices $M$.
C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
The algorithm of Appendix B transforms a CUR decomposition into a top SVD. The next algorithm makes converse transition. It uses $(m+n+r)r$ scalars and order of $(m+n)r^{2}$ flops to compute CUR decomposition having an $r\times r$ generator for a matrix given with its $r$-top SVD, for $r\leq n\leq m$. By combining this algorithm with that of Appendix B we can transform an LRA of a matrix into its CUR LRA at a low computational cost.
Algorithm C.1.
[From $r$-top SVD to CNR decomposition.] [^13] We denote nucleus by $N$ rather than $U$ or $G$ to avoid conflicts with the factor $U$ in SVD and generator $G$ of CUR of Sec. C.1.
Input:
Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{R}^{m\times n}$, $\Sigma\in\mathbb{R}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{R}^{m\times r}$, and $V\in\mathbb{R}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
Output:
Three matrices $C\in\mathbb{R}^{m\times r}$, $N\in\mathbb{R}^{r\times r}$, and $R\in\mathbb{R}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
$M=CNR.$
Computations:
By applying to the matrices $U$ and $V$ the algorithms of [14] or [33] compute two nonsingular $r\times r$ submatrices $U_{\mathcal{I},:}$ and $V^{*}_{:,\mathcal{J}}$.
Compute and output the CUR factors $C=U\Sigma V^{*}_{:,\mathcal{J}}$, $R=U_{\mathcal{I},:}\Sigma V^{*}$, and $N:=(V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1}$.
Correctness verification. Substitute the expressions for $C$, $N$ and $R$ and obtain $CNR=(U\Sigma V^{*}_{:,\mathcal{J}})((V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1})(U_{\mathcal{I},:}\Sigma V^{*})=U\Sigma V^{*}$.
Cost bounds. The algorithm uses $O(mr^{2})$ flops and $(m+n+r)r$ scalars; these cost bounds are sublinear where $r^{2}\ll n$ and are dominated at stage 2.
Let us also estimate the norm of the nucleus $||N||$. Observe that $\operatorname{rank}(V_{:,\mathcal{J}})=\operatorname{rank}(\Sigma)=\operatorname{rank}(U_{\mathcal{I},:})=r$ and $||N||\leq||V_{:,\mathcal{J}}^{-1}||~{}||\Sigma^{-1}||~{}||U_{\mathcal{I},:}^{-1}||$ and recall that $||\Sigma^{-1}||=||M^{+}||=1/\sigma_{r}(M)$.
Write $t_{q,s,h}^{2}:=(q-s)sh^{2}+1$, allow any choice of $h>1$, say, $h=1.1$, and then recall that $||U_{\mathcal{I},:}^{+}||\leq t_{m,k,h}^{a}$, $||(V_{:,\mathcal{J}}^{*})^{+}||\leq t_{n,\ell,h}^{a}$, and consequently
$||N||\leq t_{m,r,h}^{a}t_{n,r,h}^{a}/\sigma_{r}(M)$
where $a=1$ if we apply the algorithms of [14] at stage and $a=2$ if we apply those of [33].
Appendix D Generation of abridged SRHT matrices
In this section we specify the family of abridged SRHT matrices, used in Alg. 1.1. They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [26, Sec. 3.1]). The $n\times n$ matrices $H_{n}$ are obtained in $t=\log_{2}(n)$ recursive steps, but we only perform $d\ll t$ steps, and the resulting abridged matrix $H_{d,d}$ can be multiplied by a vector by using $2dn$ additions and subtractions. SRHT matrices are obtained from the matrices $H_{n}$ by means of random sampling and scaling, which we also apply to the $d$-abridged Hadamard transform matrices $H_{d,d}$. They turn into $H_{n}$ for $d=t$ but are sparse fo $d\ll t$. Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (D.1)
For any fixed pair of $d$ and $i$, each of the matrices $H_{d,i}$ is orthogonal up to scaling and has $2^{d}$ nonzero entries in every row and column; we can compute the product $MH$ for an $n\times k$ submatrix $H$ of $H_{d,d}$ by using less than $km2^{d}$ additions and subtractions.
Now define the $d$-Abridged Scaled and Permuted Hadamard matrices, $PDH_{d,d}$, where $P$ is a random sampling matrix and $D$ is the matrix of random integer diagonal scaling. Each random permutation or scaling contributes up to $n$ random parameters. We can involve more random parameters by applying random permutation and scaling also to some or all intermediate matrices $H_{d,i}$ for $i=0,1,\dots,d$.
The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf. [20, Sec. 11]).
In this Appendix we follow [40, 41, 43, 42] where abridged SRFT matrices have been defined and tested similarly to abridged SRHT matrices with similar test results.
Appendix E Cross-Approximation (C-A) iterations
In this appendix we briefly recall C-A iterations, which can be viewed as a specialization to LRA of Alternative Directions Implicit (ADI) method.
C-A iterations:
  • For an $m\times n$ matrix $M$ and a target rank $r$, fix four integers $k$, $\ell$, $p$ and $q$ such that $0<r\leq k\leq p\leq m$ and $r\leq\ell\leq q\leq n$. [C-A iterations are simplified if $p:=k$ and $q:=\ell$.]
  • Fix an $m\times q$ “vertical” submatrix of the matrix $M$, made up of its $q$ columns. [^14] Of course, one can alternatively begin C-A iterations with a “horizontal” submatrix.
  • By applying a fixed CUR LRA subalgorithm, e.g., one of the algorithms of [29, 14, 33, 8], [^15] Such a subalgorithm runs at sublinear cost on the inputs of smaller size involved in C-A iterations, even where the original algorithms of [29, 14, 33, 8] applied to an $m\times n$ matrix $M$ run at superlinear cost. compute a $k\times\ell$ CUR generator $G$ of this submatrix and reuse it for the matrix $M$.
  • Compute the rank-$r$ CUR of $M$ based on the generator $G$ and output this CUR it if it is close enough.
  • Otherwise swap $p$ and $q$ and reapply the algorithm to the matrix $M^{*}$.
    [This is equivalent to computing a $k\times\ell$ CUR generator of a fixed $p\times n$ “horizontal” submatrix $M_{1}$ of $M$ that covers the submatrix $G$.]
  • Recursively alternate such “vertical” and “horizontal” steps until an accurate CUR LRA is computed or until the number of recursive C-A steps exceeds a fixed tolerance bound.
Figure 2: The first three recursive C-A steps output three striped submatrices.
Initialization recipes. For a large class of input matrices it is efficient to apply adaptive C-A iterations (cf. [1, 2]). They combine initialization with dynamic search for gaps in the singular values of $M$ by properly adapting Gaussian elimination with pivoting. The alternative initialization in [16] uses $O(nr^{2})$ flops to initialize the C-A iterations for an $n\times n$ input and $q=s=k=\ell=r$, and then the algorithm uses $O(rn)$ flops per C-A step. The output of C-A iterations, however, can be off by a significant factor.
Appendix F Iterative refinement of LRA with no compression
The definition of $r$-truncation immediately implies
Lemma F.1.
$M_{s}+(M-M_{s})_{r}=M_{r+s}$ for any pair of positive integers $r$ and $s$.
Next, based on this lemma, we lift $r$-top SVD into $kr$-top SVD by applying iterative refinement of LRA with no compression.
Algorithm F.1.
[Lifting $r$-top SVD into $kr$-top SVD.]
Input:
An $m\times n$ matrix $M$, $n\leq m$, two positive integers $k$ and $r$, an Alg. $\mathbb{S}$ that adds or subtracts a pair of $m\times n$ matrices, and an Alg. $\mathbb{T}$ that computes the $r$-truncation of an $m\times n$ matrix.
Output:
The $kr$-truncation of $M$.
Initialization:
Write $X_{0}:=0$ and $E_{0}:=M$.
Computations:
Recursively combine Algs. $\mathbb{S}$ and $\mathbb{T}$ to compute the matrices
$X_{i+1}:=X_{i}+(E_{i})_{r},~{}E_{i+1}:=M-X_{i+1},~{}~{}i=0,1,\dots,k-1.$ (F.1)
Output the matrix $X_{k}$.
The following proposition implies correctness of the algorithm.
Proposition F.1.
It holds that $X_{i}=M_{ir}$ for the matrices M and $X_{i}$ of Alg. F.1 for $i=0,1,\dots,k$.
Proof.
The claim holds for $i=0$ because $M_{0}=0$. Let it hold for $i\leq j$, that is, let $X_{j}=M_{jr}$, and then deduce from Eqn. (F.1) for $i=j$ that $X_{j+1}=X_{j}+(M-X_{j})_{r}$. By assumption, $X_{j}=M_{jr}$. Combine the latter equations and obtain $X_{j+1}=M_{jr}+(M-M_{jr})_{r}$. Apply Lemma F.1 for $s=jr$ and obtain that $X_{j+1}=M_{(j+1)r}$. Now the proposition follows by induction. ∎
Trying to extend our error analysis of Alg. 3.1 to Alg. F.1 we needed to estimate for the norm of the perturbation of the $r$-top SVD, but at this point the known studies (cf. [15, Thm. 8.6.5], [52, Thm. 3.3.4]) provided much weaker results than Thm. 1.2, and we ended with no support for numerical stability of Alg. F.1.

### v16

@@
-Accurate Low Rank Approximation at a Low Computational Cost
+Near-Optimal Very Low Rank Approximation at Sublinear Cost
 Abstract
-Given an algorithm $\mathbb{A}$ for computing a crude Low Rank Approximation (LRA) of a matrix having fast decaying spectra of singular values, we greatly decreased the spectral norm of its output error matrix at sublinear additional computational cost.
-Furthermore, for a variety of synthetic and real world inputs we tested numerically variants where our entire algorithms run at sublinear cost, and in most cases we still obtained accurate or even near-optimal LRAs.
-Towards enhancing our progress to a larger class of matrices, we nontrivially extended to LRA the popular technique of iterative refinement.
+Eckart-Young’s theorem implies that even optimal rank-$r$ approximation of an $m\times n$ matrix $M$ for $m\geq n$ is only interesting if there is a gap in the segment of the $r+1$ largest singular values of $M$.
+If $M$ also has a gap in the segment of the next $p$ largest singular values, then one can compute its rank-$r$ approximation with random sampling and oversampling in nearly linear arithmetic time.
+For matrices $M$ having such a double gap in their spectra and for $n$ greatly exceeding $(r+p)^{2}$, our algorithm runs in sublinear time, and for a large subclass of the class of such matrices $M$ its output rank-$r$ approximation of $M$ is near-optimal under the spectral norm.
+We also extend the classical iterative refinement to LRA and analyze and test the resulting algorithm.
 Key Words:
-low rank approximation, near-optimal algorithms, sublinear cost algorithms, singular values, gaps in the spectrum, iterative refinement
+low rank approximation, singular values, gaps in the spectrum, sublinear cost algorithms, iterative refinement
 2020 Math. Subject Classification:
 65F55, 65N75, 65Y20, 68Q25, 68W20
-1 LRA of a matrix: State of the Art and our Novelties
+1 LRA of a matrix: the State of the Art and our Novelties
+Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra.
+We first briefly recall the State of the Art, referring to [11, 33, 18] for broader expositions, and then outline our LRA algorithms.
 1.1 The State of the Art
-Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra.
-[^1] The terms “low”, “large”, “small”, “close”, “near” etc are used in context in our introductory comments.
-Next we outline its State of the Art, referring for broader expositions to foremost papers [20, 54].
 1.1.1.
-Some assumptions.
-For simplicity we seek LRA just in the field $\mathbb{R}$ of real numbers but can readily extend our studies to the complex field $\mathbb{C}$.
-We assume that an $m\times n$ input matrix $M$ is a black box matrix, given by an oracle (black box subroutine) for computing its product with a vector rather than by the entries.
-This is a natural model, e.g., for the study of matrices of immense sizes coming from Big Data, for which the access to all matrix entries is too costly (see further motivations in [54, Sec.
+LRA and CUR LRA.
+Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors.
+[^1] The sketch of a matrix in [36, 33] is its product with a fixed set of vectors.
+Then we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [33, Sec.
 1.3]).
-Under this model linear operations with matrix-by-vector products can be performed with vectors.
-Hereafter $a\ll b$ means that $a=o(b)$ but also that $b$ greatly exceeds $a$ in non-asymptotic context, $A^{+}$ denotes the Moore-Penrose pseudo inverse of $A$, and $m\geq n$; otherwise we can work with the $n\times m$ transpose $M^{T}$.
-[…]
-Computation of $M_{r}$ and $r$-top SVD involves order of $mn^{2}$ floating point operations (flops) and the memory space for about $mn$ scalars (see [15, Fig.
-8.6.1]), although the output $r$-top SVD only uses memory for $(m+n+r)r$ scalars.
+Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [35], that is, admit LRA, with which one can operate superfast, aka at sublinear cost – by using much fewer than $mn$ scalars and flops.
+[^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [24], and we extend this concept to other matrices.
+Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r\ll n$ or a fixed target range [^3] Here and hereafter $a\ll b$ means that $b$ greatly exceeds $a$ in context, e.g., $a=o(b)$ as $b\mapsto\infty$.
+for such a rank (actually, a proper choice of such targets is a challenge [11, Sec.
+4.2]) and then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
+$X\in\mathbb{C}^{m\times k},Y\in\mathbb{C}^{k\times\ell},~{}Z\in\mathbb{C}^{\ell\times n},$ (1.1)
+$k\ll m,~{}\ell\ll n,~{}\min\{k,\ell\}=r,$ (1.2)
+and the spectral and/or Frobenius norm of the matrix $M-XYZ$ is small in context, say, close to optimal.
+The matrices $X,Y,Z$ have $mk+k\ell+\ell n\ll mn$ entries overall, and we can multiply the matrix $XYZ$ by a vector by using less than $2(mk+k\ell+\ell n)\ll mn$ flops.
+If the matrices $X$ and $Z$ are made up of the columns and rows of $M$, that is, are defined by means of pre- and post-multiplication of $M$ by selected coordinate vectors, respectively, then write $X:=C$, $Z:=R$, and $Y:=U$ and call the triplet $\{X,Y,Z\}$ a CUR LRA.
+1.1.2.
+LRA from top SVD, SVD of an LRA, and CUR from SVD.
+By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, obtain a matrix $M_{r}$ of rank $r$, said to be $r$-truncation of $M$.
+By virtue of the Eckart-Young-Mirsky theorem [7, Thm.
+2.4.8], $M_{r}$ is an optimal rank-$r$ approximation444“Rank- $r$ approximation” means “rank at most $r$ approximation” throughout.
+of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
+Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [7, Fig.
+8.6.1]).
+This is too expensive for computing LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ and then its CUR LRA by using $(m+k)k+(n+\ell)\ell$ scalars, $O(mk^{2}+n\ell^{2})$ flops and $(m+n+r)r$ scalars, $O(mr^{2})$ flops, respectively (see Appendices B and C.2, respectively).
 1.1.3.
-Variations and computational equivalence of LRA tasks.
-Next we state LRA task by generalizing this and other known representations of LRA as follows: given a matrix $M\in\mathbb{R}^{m\times n}$ and a pair of integers $r_{-}$ and $r_{+}$ such that
-$0<r_{-}\leq r_{+}\ll n\leq m,$ (1.1)
-fix two integers $k$ and $\ell$ such that
-$k\ll m,~{}\ell\ll n,~{}r\leq\min\{k,\ell\},~{}{\rm and}~{}r_{-}\leq r\leq r_{+}$ (1.2)
-and seek a rank-$r$ approximation of $M$ by a triplet of matrices $X\in\mathbb{R}^{m\times k}$, $Y\in\mathbb{R}^{k\times\ell}$, and $Z\in\mathbb{R}^{\ell\times n}$ such that the error matrix $M-XYZ$ has a small spectral and/or Frobenius norm.
-[^3] For $r_{-}=r_{+}$, relationships (1.2) define a target rank $r=r_{-}=r_{+}$.
-Otherwise, more realistically, they define a target range for rank $r$.
-If the matrices $X=C$ and $Z=R$ are made up of the columns and rows of $M$, respectively, then $Y$ is commonly denoted by $U$ (or by $G$ by the followers of the pioneering work [17]) and the triplet $\{X,Y,Z\}$ is called CUR LRA (or CGR LRA) and is a particularly memory efficient LRA.
-For any $r\leq\min\{k,\ell\}$ we can transform any LRA triplet $\{X,Y,Z\}$ for $M$ into its $r$-top SVD and further into its CUR (CGR) LRA by using $O(mk^{2}+n\ell^{2})$ flops, $(m+k)k+(n+\ell)\ell$ scalars and $O(mr^{2})$ flops, $(m+n+r)r$ scalars, respectively (see Appendices B and C.2, respectively).
-1.1.4. LRA via random sampling. Impressive progress in LRA in the last decades relied on random sampling algorithms (see [8, 6, 26, 20, 56], and the references therein).
-We need an efficient LRA algorithm for incorporation into our Alg.
+LRA via random sampling and oversampling.
+Random sampling with oversampling enables fast computation of LRA (see [3, 17, 11, 36, 33, 20], and the references therein).
+For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [11] successively generates a random test matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
+Remark 1.1.
+“In practice, the target rank $k$ is rarely known in advance.
+Randomized algorithms are usually implemented in anadaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [11, Sec.
+4.2]).
+In the best studied case the test matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables.
+In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [11, Thm.
+10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops.
+The parameter $p$ and the size of the test matrix $H$ explode as $\epsilon\mapsto 0$, but LRA with $\epsilon$ as large as 1/2 or 1 is still valuable.
+In other LRA algorithms (e.g., in [4, 3, 33, 20]) $M$ is both pre- and post-multiplied by random test matrices $F$ and $H$, respectively.
+In particular, the TYUC algorithm of [33], modifying those of [3, Thm.
+4.7] and [3, Thm.
+4.3, display 1], and the generalized Nyström approximation GN of [20] first compute the products $FM$ and $MH$ for random test matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$ with the expected Frobenius error norm within a factor of 2 from optimal in the case of random Gaussian test matrices $F$ and $H$ (namely, the norm does not exceed the optimal one for rank-$\frac{k}{2}$ approximation).
+For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [20]) and $O(\ell mn)$ flops.
+The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [11, Secs. 10 and 11], [36, Ch. 6.2], [5, Thm. 3.4]).
+1.1.4.
+Superfast LRA for special subclasses of matrices.
+For some special but large and important classes of matrices, most notably for Symmetric Positive Semidefinite (SPSD) matrices but also for Distance matrices and a subclass of Cauchy matrices, the superfast randomized algorithms of [19, 1, 12, 14, 2] compute LRA whose expected Frobenius error norm is compared well with the optimal bound.
+The mere existence of such LRA algorithms is surprising for $n\times n$ SPSD matrices, whose class includes matrices $A^{*}A$, defined by nonsingular $n\times n$ matrices $A$ with $n^{2}$ independent entries.
+Notice that the superfast algorithm of [15, Part III] for SPSD matrices is deterministic, but its spectral and Frobenius output error norms exceed their optimal values by at most a factor of $n$.
+Unfortunately, one cannot compute LRA superfast for every matrix: already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [26, 27, 28]).
+If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0.
+1.2 Our LRA algorithm
+Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$.
+We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$ of size $m\times n$ such that $r<\rho\ll\sqrt{n}$.
+Our algorithm is superfast and, for a large subclass of $(r,\rho)$-matrices $M$, its output very low rank approximation (VLRA) is near-optimal under spectral norm.
+We rely on combining three surprisingly simple observations, namely, that (i) random sampling with oversampling is the only stage where HMT, TYUC, and GN algorithms are not superfast for VLRA; (ii) we can perform this stage superfast by using sparse test matrices (in particular abridged SRHT matrices of Appendix D), and (iii) while the accuracy of the VLRA output in step (ii) tends to be unsatisfactory, we can readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$.
+Our LRA Alg.
 2.1 of Sec.
-2 and for natural candidates consider random sampling algorithms, such as the following algorithm of [54], which modifies those of [6, Thm.
-4.7] and [6, Thm.
-4.3, display 1].
-Algorithm 1.1.
-[LRA by means of random sampling].
-Input:
-An $m\times n$ black box matrix $M$, given implicitly by a subroutine for its multiplication by vectors, and positive integers $r_{-}$ and $r_{+}$ satisfying (1.1).
-Output:
-An integer $r$ satisfying (1.2) and, moreover, equal to $\min\{k,\ell\}$ and a triplet of matrices $\{X,Y,Z\}$ such that $XYZ$ is a rank-$r$ matrix close to $M$.
-Initialization:
-Fix a pair of integers $k$ and $\ell$ satisfying (1.2) and generate a pair of independent random test matrices $F\in\mathbb{R}^{\ell\times m}$ and $H\in\mathbb{R}^{n\times k}$.
-Computations:
-  1. 1.
-    Compute the matrices $Z:=FM\in\mathbb{R}^{\ell\times n}$ and $MH\in\mathbb{R}^{m\times k}$, which are two sketches (samples) of $M$.
-  2. 2.
-    Compute the matrix $X\in\mathbb{R}^{m\times k}$ denoting the Q factor of the thin QR factorization of $MH$ – (cf. [15, Thm. 5.3.2]).
-  3. 3.
-    Compute the matrx $Y:=(FQ)^{+}\in\mathbb{R}^{k\times\ell}$.
-  4. 4.
-    Output the triplet of matrices $\{X,Y,Z\}$.
-The algorithm involves $O((k+\ell)mn)$ flops [^4] One can do with just $O(mn\log(k+\ell))$ flops to ensure a little weaker bounds on failure probability [20, Sec.
-11.2].
-and about $(m+n)(k+\ell)$ scalars, not counting the cost of the generation of random matrices $F$ and $H$.
-Like various other random sampling LRA algorithms, Alg.
-1.1 is universal, that is, devised for all matrices $M\in\mathbb{R}^{m\times n}$; for proper pairs of dense test matrices $f$ and $H$ and integers $k$ and $\ell$ in $O(r\log(r))$ it belongs to the class of relative error LRA algorithms: for any fixed pair of positive $\epsilon$ and $\delta$ it outputs rank-$r$ approximations of $M$ whose Frobenius error norm is within a factor of $1+\epsilon$ from the optimal value $||M-M_{r}||_{F}$ with a failure probability at most $\delta$ (see the bibliography cited above).
-1.1.5.
-Superfast (sublinear cost) LRA.
-The LRA algorithms of [28, 4, 21, 23, 5] (randomized) and [24, Part III] (deterministic) are known to run superfast, aka at sublinear cost, that is, to use $o(mn)$ scalars and flops for a fixed pair of $\epsilon$ and $\delta$ [^5] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [34, 36], and we extend this concept to other matrices.
-Unfortunately, already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [40, 41, 42, 43]).
-If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0.
-Likewise, any deterministic or randomized superfast estimation of the error norm of an LRA fails on the matrices of that small family.
-Hence a superfast algorithm cannot be universal for computation of LRA or even for its a posteriori error norm estimation.
-Accordingly, the randomized LRA algorithms known to be superfast are restricted to some large and important classes of input matrices such as Symmetric Positive Semidefinite (SPSD) matrices, Distance matrices, or a subclass of Cauchy matrices.
-Furthermore, strictly speaking they are NOT superfast because they involve a posteriori error norm estimation in linear or superlinear time.
-The latter deficiency is only avoided in the deterministic algorithm of [24, Part III] for LRA of SPSD matrices – its output error norm is estimated as by-product, at no additional cost.
-1.1.6.
-Conflicts #$\#1$ for LRA.
-Trying to optimize simultaneously various permissible parameters of LRA such as arithmetic cost and the costs of randomization, storage, and communication, representing the computational cost one runs into the conflict with the quality of a priori error bounds one (see [54, Sec.
-3.9]), which boils down to the conflict between the two main goals of devising efficient LRA algorithms, that is, minimization of (a) the output error norm and (b) represented by the memory space and number of flops involved.
-Here are some specific samples of this omnipresent LRA conflict #$\#1$, routinely resolved by means of trade-off.
-(i) The known random sampling LRA algorithms need unbounded memory size and flop number to ensure with a high probability (whp) the relative error norm bound within $1+\epsilon$ from optimal under the Frobenius norm for $\epsilon\rightarrow 0$ [20, Sec.
-10.3].
-Furthermore, the LRA algorithms supporting the same output error bound under the spectral rather than Frobenius norm are more costly (see [56, Ch.
-6.2], [13, Thm.
-3.4]).
-(ii) Adaptive Cross-Approximation (ACA) iterations, [^6] We briefly recall their simplified version of Cross-Approximation (C-A) iterations in Appendix E.
-which are empirically superfast for a large class of matrices $M$, output LPAs whose error norm bound is at best within a factor of $2\sqrt{mn}$ from optimal under both spectral and Frobenius matrix norms (see [30, Cor.
-2]).
-(iii) Likewise, the output error norms of the relative error LRAs of [24, Part III] for $n\times n$ Symmetric Positive Semidefinite (SPSD) matrices $M$ may exceed the optimal spectral and Frobenius norms $||M-M_{r}||$ by a factor of $n$.
-1.2 Our progress (briefly)
-For matrices with fast decaying spectra of singular values (hereafter FDS matrices) our algorithms decrease the spectral norm of the output error matrix of the known LRA algorithms at sublinear additional cost, that is, superfast.
-Our numerical tests with both synthetic and real world matrices are in good accordance with these formal estimates.
-Furthermore, some variants of our entire algorithms run superfast, and we empirically observe that they still output accurate or even near-optimal LRAs for a large subclass of FDS matrices.
-1.3 LRA by means of escalation
-Next we outline our rather simple techniques.
-1.3.1.
-More definitions and two theorems.
-Hereafter $\sigma_{j}(M)$ denotes the $j$th top (largest) singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$.
-$||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Sec.
+2 relies on these observations.
+We compute at first (1) a crude but reasonably close rank-$\rho$ approximation $M(\rho)$ of $M$ and then (2) the rank-$r$ truncation $M(\rho)_{r}$ of $M(\rho)$, which we output as rank-$r$ approximation of $M$.
+LRA via random sampling with oversampling fits this outline, but based on above observations (i)–(iii) we run the algorithm superfast for $\rho\ll\sqrt{n}$ and argue that its output VLRA is near-optimal for a large subclass of $(r,\rho)$-matrices.
+We can perform stage (1) of the algorithm superfast by applying any of the HMT, TYUC and GN algorithms for $k:=\rho\ll\sqrt{n}$ with sparse test matrices such that their multiplication by $M$ is superfast (see [13], [33, Sec.
+3.9], and our Sec.
+3).
+Apart from the stage of multiplication by test matrices, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$, and obtain Moore-Penrose pseudoinverses of a $k\times\ell$ matrices ${\bf\Psi Q}$ in [33] and $R_{\epsilon}$ in [20] by using $O(k^{2}\ell)$ flops (see [7, Fig.
+8.6.1]).
+These computations are superfast where $(k+p)^{2}\ll n$ and $k^{2}\ll n$, respectively, for $\ell\leq 2k$, say.
+At Stage (2) we apply Alg.
+B.1.
+It computes SVD and $r$-truncation of $M(\rho)$ by using about $(m+n)\rho$ scalars and $O(mr^{2})$ flops, that is, $\ll m\sqrt{n}$ scalars and $\ll mn$ flops for $\rho\ll\sqrt{n}$.
+By virtue of Thm. 2.3, the spectral norm $||\cdot||:=||\cdot||_{2}$ of the output error matrix of Alg. 2.1 satisfies
+$||M-M(\rho)_{r}||\leq||M-M_{r}||+2||M-M(\rho)||,$ (1.3)
+that is, Alg.
+2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$.
+For $(r,\rho)$-matrices, this factor can be large only if the superfast computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails.
+It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse test matrices tend to be unsatisfactory as LRA [13], [33, Sec.
+3.9], but in our numerical tests in Sec.
+5 accuracy of the output matrices of stage (1) of Alg.
+2.1 tended to be raised to near-optimal level at stage (2).
+1.3 Iterative LRA
+In Sec.
+4 we apply our LRA algorithm to the output error matrix, hoping to decrease its norm if it is too large.
+Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs.
+The algorithm is essentially reduced to recursive applications of Alg.
+2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance.
+We counter this deficiency by recursively compressing the sums into rank-$r$ matrices, although this limits output accuracy of computed LRAs.
+The resulting iterative LRA algorithm is a natural extension to LRA of an efficient popular technique of matrix computations; we specify, analyze, and test it in Secs.
+4 and 5.
+We devised this algorithm to complement and strengthen Alg.
+2.1, but it is essentially reduced to recursive applications of Alg.
+2.1.
+Our upper estimates for the accuracy of both algorithms are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
+In our tests, the output accuracy of our iterative LRA applied to $(r,\rho)$-matrices improved significantly at the second iteration, little improved afterwards, and overall error norm was never less than with Alg.
+2.1 for the upper rank $\rho=4r$.
+1.4 Related works
+Our study was prompted by extensive work on LRA in the last decades (see [11, 33, 18] and the bibliography therein) and in particular by fast LRA based on random sampling and oversampling.
+To accelerate these algorithms we use abridged SRHT matrices studied in the papers [26, 27, 28], which in turn extends the earlier study of Gaussian elimination without pivoting in [29, 31].
+Our iterative LRA extends the algorithms of [16, 25], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [32, Secs.
+3.3.4 and 4.2.5], [10, Ch.
+12 and Sec.
+20.5], [7, Secs.
+3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [32, page 223 and the references on page 225]).
+Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [22, 30].
+Actually, we first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.
+1.5 Organization of our paper
+We present our accelerated oversampling LRA algorithm in Sec.
+2, discuss superfast computation of a crude LRA in Sec.
+3, present our iterative LRA algorithm in Sec.
+4, and cover our numerical experiments in Sec.
 5.
-In particular we only recall the Eckart-Young-Mirsky theorem (cf. [50, 12, 27], [15, Thm. 2.4.8]) under the spectral norm:
+We devote our short Sec.
+6 to conclusions.
+In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails.
+In Appendix B we compute $r$-top SVD of an LRA of a low-rank matrix superfast.
+In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA.
+In Appendix D we generate abridged SRHT matrices, whose incorporation as test matrices makes random sampling LRA superfast.
+2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
+We begin with some background materials.
+Hereafter $||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Remark 2.2.
+In particular we only use Eckart-Young’s (rather than Mirski’s) theorem:
 Theorem 2.1.
@@
 Theorem 2.2.
-[15, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
+[7, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
 $|\sigma_{j}(M+E)-\sigma_{j}(M)|\leq||E||~{}{\rm for}~{}j=1,\dots,\min\{m,n\}.$
-1.3.2.
-Gaps in the spectrum and escalation of LRAs.
-For a fixed $r$ in a target range, we fix an upper rank $\rho$ satisfying $r\leq\rho\ll n$ and then (i) compute a rank-$\rho$ approximation of $M$ given by a triplet $\{X,Y,Z\}$ such that
-$||M-M(\rho)||=\beta\sigma_{\rho+1}(M),~{}{\rm for}~{}M(\rho)=XYZ,~{}1\leq\beta\leq\beta_{+}~{}{\rm and~{}a~{}fixed~{}bound}~{}\beta_{+},$ (1.3)
-and (ii) compute and output the $r$-top SVD of $M(\rho)$.
-The relative error random sampling algorithms of [20, 54] use from $O(mn\log(\rho))$ to $O(mn\rho)$ flops and about $(m+n)\rho$ scalars to perform stage (i), while their heuristic variations in our Sec.
-4 perform stage (i) at a lower cost; deterministic Alg.
-B.1 of Appendix B uses $O(m\rho^{2})$ flops and about $(m+n)\rho$ scalars to perform stage (ii).
-[^7] By applying the algorithm of [14] instead of Alg.
-B.1 at stage (ii), one can a little accelerate the computations while only slightly increasing the error norm bound.
-In Alg. 2.1 of Sec. 2 we elaborate upon this outline and deduce that
-$||M-(M(\rho))_{r}||\leq\sigma_{r+1}(M)+2\beta\sigma_{\rho+1}(M)=\Big{(}1+\frac{2\beta\sigma_{\rho+1}(M)}{\sigma_{r+1}(M)}\Big{)}\sigma_{r+1}(M).$ (1.4)
-The output error norm $||M-M(\rho)_{r}||$ of Alg.
-2.1 is small if the value $\frac{\sigma_{r+1}(M)}{2\beta\sigma_{\rho+1}(M)}$ is large, that is, if there is a large gap in the segment $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of the spectrum of $M$.
-For comparison, a rank-$r$ approximation of $M$ is only interesting if the value $\frac{\sigma_{1}(M)}{\sigma_{r+1}(M)}$ is large, that is, if there is a gap in the leading segment $[\sigma_{1}(M),\sigma_{r+1}(M)]$ of the spectrum.
-We can reveal gaps in the spectrum by approximating the values of $||M-(M(\rho))_{r}||$ for various pairs of $\rho$ and $r$ and comparing the output error norms.
-1.3.3.
-Superfast variations.
-We can compute the $r$-truncation of $M(\rho)$ superfast for any matrix $M$ if $\rho$ is small such that $\rho^{2}=o(n)$, while some superfast variations of the known LRA algorithms (e.g., those of [20, 54] discussed in our in Sec.
-4) output reasonably accurate rank-$\rho$ approximations $M(\rho)$ for a large class of FDS matrices $M$.
-We cannot estimate the spectral error norm of a matrix superfast, and even its LAPACK heuristic estimation is not superfast, but we propose a heuristic superfast 1-norm estimator in Sec.
-5 and test it in Sec.
-6.4.
-1.3.4.
-Iterative LRAs.
-In Alg.
-3.1 of Sec.
-3, which is a competitive alternative to Alg.
-2.1, we write $X_{0}:=0$, $E_{i}:=M-X_{i}$, and $X_{i}:=(X_{i-1}+E_{i}(\rho))_{r}$ for $i=1,2,\dots,k$.
-We compress the matrices $X_{i-1}+E_{i}(\rho)$, that is, keep their ranks bounded by $r$ – by working with their $r$-truncation, and we do this superfast as long as their rank bound $\rho+r$ is low.
-For computation of the matrix $E_{i}(\rho)$ we use the algorithms of stage (i) of Alg.
-2.1.
-This way we extend to LRA the popular iterative refinement algorithms.
-The algorithm is essentially reduced to $k$ applications of Alg. 2.1.
-In our tests Alg.
-3.1 output LRAs with relative output errors about 1.1 for a matrix class marked Shaw for which Alg.
-2.1 failed, but for all other inputs Alg.
-3.1 was less accurate than Alg.
-2.1 applied with the upper rank $\rho$ twice as large versus Alg.
-2.1.
-Our upper estimates for the accuracy of both algorithms are close to one anaother but are both pessimistic, and it seems that further comparative study of the two algorithms is in order.
-For completeness, we cover iterative refinement of LRA with no compression in Appendix F, but unlike Alg. 3.1, we cannot provide any support for its numerical stability.
-1.4 Related works
-Motivated by exposition of the progress in LRA in [20, 54], we extend the rudimentary study of LRA escalation in [40, 41, 43, 42] (cf. the end of Sec. 4).
-The algorithms in [25, 39] extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [51, Secs.
-3.3.4 and 4.2.5], [19, Ch.
-12 and Sec.
-20.5], [7], [15, Secs.
-3.5.3 and 5.3.8], [3, Secs.
-1.4.6, 2.1.4, and 2.3.7]) but also widely applied in various other linear and nonlinear matrix computations (see [51, page 223 and the references on page 225]).
-Its extension to LRA is not straightforward – because the rank of the computed approximations grows recursively; we block this growth by applying re-compression, tracing such techniques to [31, 32, 37, 46, 38, 35] and [34, Secs.
-6.3-6.5].
-Very different iterative refinement of LRA in [25] uses no crude auxiliary LRA but instead bounds the Principal Angle Distances between singular spaces of $M$ and its LRAs and incorporates randomization based on the leverage scores of [8].
-Historically, we first succeeded with iterative refinement of LRA and then noticed that escalation alone is quite competitive.
-In our choice of abridged SRHT matrices for sparse multipliers we follow [40, 41, 42, 43], which proved some crude but reasonable probabilistic upper bounds on output errors of some superfast LRA algorithms under randomization in the input matrix space defined by a fixed probability structure in that space.
-[^8] These papers continued the study in [45, 44, 47], which introduced a probability structure in the space of input matrices to motivate application of Gaussian elimination without pivoting.
-The relevance of the gaps in the spectrum of the singular values of a matrix to its LRA is well-known but has not been exploited in the way that we propose, test, and analyze.
-Likewise, from the comments of an expert on an earlier draft of this paper, we learned that [9] also worked with error matrices of LRA, but in the directions quite different from ours.
-1.5 Organization of our paper
-We present our LRA escalators in Secs.
-2 and 3 and discuss superfast computation of a crude LRA in Sec.
-4.
-We specify a superfast heuristic 1-norm estimator in Sec.
-5, cover our tests in Sec.
-6, and devote our brief Sec.
-7 to conclusions.
-In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails.
-In Appendix B we compute $r$-top SVD of an LRA or any low-rank matrix superfast.
-In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA.
-In Appendix D we generate abridged SRHT matrices.
-In Appendix E we briefly recall ACA (C-A) iterations.
-In Appendix F we iteratively refine LRA with no compression.
-2 LRA by means of escalation
-Next we elaborate upon escalation outlined in Sec. 1.3.2.
+In the following algorithm we elaborate upon the outline of Sec. 1.2 assuming that we are given two black box subroutines that output a crude LRA of a matrix.
 Algorithm 2.1.
-LRA by means of escalation.
+LRA of $(r,\rho)$-matrices.
 Input:
-Two positive integers $m$ and $n\leq m$, an $m\times n$ matrix $M$, possibly given implicitly by its sketches, three positive values $\widehat{\tau}$, $\widehat{\mu}$, and $r$, $r\ll n$, denoting bound on the overall computational cost, [^9] Here and hereafter we use a single character $\tau$ to represent both flops and memory size involved.
-an absolute error norm bound, and a fixed integer in a target range, respectively; three auxiliary Algs.
-$\mathbb{A}$, $\mathbb{B}$, and $\mathbb{C}$ which for any $\rho$, $r\leq\rho\leq n$, compute a (crude but reasonably close) rank-$\rho$ approximation of $M$, compute $r$-top SVD of a rank-$\rho$ matrix of size $m\times n$, and estimate the norm of the difference of $M$ and another $m\times n$ matrix, respectively, by performing within computational cost bounds $\tau(\rho)$, $\tau(\rho,r)$, and $\tau^{\prime}$, respectively.
-[We elaborate upon Algs.
-${\bf\mathbb{A}}$, ${\bf\mathbb{C}}$, and ${\bf\mathbb{B}}$ in Secs.
-4 and 5 and Appendix B, respectively.]
+An $m\times n$ matrix $M$, for $n\leq m$, given implicitly by an oracle (black box subroutine) for its multiplication by a vector (cf.
+Sec.
+1.1.1), a target rank $r\ll n$, and a tolerance $\mu$ to the output error norm.
 Output:
-FAILURE or SVD of a rank-$r$ matrix $X\approx M$ and the norm $||M-X||$ not exceeding $\widehat{\mu}$.
+FAILURE or SVD of a rank-$r$ matrix $X\approx M$ such that $||M-X||\leq\mu$.
 Initialization:
-Fix the maximal upper rank $\rho$ satisfying the following bounds (see Remark 2.1),
-$\tau(\rho)+\tau(\rho,r)+\tau^{\prime}\leq\widehat{\tau}~{}{\rm and}~{}r\leq\rho\ll n\leq m.$ (2.1)
+Fix an upper rank $\rho$ satisfying
+$r\leq\rho\ll\sqrt{n}.$ (2.1)
 Computations:
   1. 1.
-    Apply Alg. $\mathbb{A}$ to output a rank-$\rho$ approximation of $M$ given by SVD of some matrix $M(\rho)$.
+    Compute an approximation of $M$ given by SVD of a matrix $M(\rho)$ of rank $\rho$.
   2. 2.
-    Apply Alg. ${\bf\mathbb{B}}$ to that matrix to compute its $r$-top SVD,101010 Alg. B.1 of Appendix B performs this stage superfast where $\rho^{2}\ll n$. which is SVD of $X:=M(\rho)_{r}$.
+    Compute its $r$-top SVD (cf. Appendix B), which is SVD of $X:=M(\rho)_{r}$.
   3. 3.
-    If Alg. ${\bf\mathbb{C}}$ shows that
-    $\mu:=||E||\leq\widehat{\mu}~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
-    then output that SVD and the estimated norm $\mu$. Otherwise output FAILURE.
-Remark 2.1.
-[The computational cost bound, the choice of an upper rank, and a simplified version of Alg.
-2.1.] (i) Bound (2.1) implies that the overall computational cost $\tau(\rho)+\tau(\rho,r)+\tau^{\prime}$ stays within tolerance $\widehat{\tau}$.
-(ii) Only limited by Eqn.
-(2.1), we maximize the upper rank $\rho$ – to decrease the error norm bound $\mu$.
-(iii) By increasing the upper bound $\widehat{\mu}$ on the output error norm $\mu$, we can decrease the cost of computations at stage 2, which, however, tends to be dominated.
-(iv) In Sec.
-6.3 we test a simplified version of Alg.
-2.1 where we impose no upper bounds $\widehat{\tau}$ and $\widehat{\mu}$ and display the output error norms for $\rho:=ir$ and $i=2,3,4,5$.
+    Estimate the norm $||E||$. If
+    $||E||\leq\mu~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
+    then output that SVD and $||E||$. Otherwise output FAILURE.
 Next we prove bound (1.3).
 Theorem 2.3.
-For $\mu$ of Eqn. (2.2) and $\beta$ of (1.3), it holds that
-$\mu\leq\sigma_{r+1}(M)+\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||=\beta\sigma_{\rho+1}(M).$ (2.3)
+The error norm $||E||=||M-X||$ (cf. Eqn. (2.2)) satisfies
+$||E||\leq\sigma_{r+1}(M)+2\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||,$ (2.3)
+which turns into (1.3) for $\sigma_{r+1}(M)=||M-M_{r}||$.
 Proof.
-Represent $E$ of (2.2) as $M-M_{r}+M_{r}-M(\rho)_{r}$, apply triangle inequality, and obtain
-$||E||\leq||M-M_{r}||+||M_{r}-M(\rho)_{r}||=\sigma_{r+1}(M)+||M_{r}-M(\rho)_{r}||.$
-Thm. 1.2 implies that $||M_{r}-M(\rho))_{r}||\leq\nu$. Combine the latter two bounds. ∎
-Corollary 2.1.
-[Relative output error norms of Algs. $\mathbb{A}$ and 2.1.] Write (cf. (1.3))
-$\alpha:=\alpha(M,\rho,r):=\frac{||E||}{\sigma_{r+1}(M)}-1~{}{\rm and}~{}\beta:=\beta(M,\rho):=\frac{\nu}{\sigma_{\rho+1}(M)}.$ (2.4)
-Then the rank-$r$ approximation of $M$ output by Alg.
-2.1, for a fixed pair of $r$ and $\rho$, is optimal up to within the absolute and relative error norms $\nu$ and $1+\alpha$, respectively, such that
-$1+\alpha\leq\frac{\nu}{\sigma_{r+1}(M)}=\frac{\beta\sigma_{\rho+1}(M)}{\sigma_{r+1}(M)}.$ (2.5)
+Represent $E$ of (2.2) as $M-M(\rho)+M(\rho)-M(\rho)_{r}$, apply triangle inequality, and obtain
+$||E||\leq||M-M(\rho)||+||M(\rho)-M(\rho)_{r}||=\nu+\sigma_{r+1}(M(\rho)).$
+Thm. 2.2 implies that $\sigma_{r+1}(M(\rho))\leq\nu+\sigma_{r+1}(M)$. Combine the latter two bounds. ∎
 Remark 2.1.
-How sharp are the upper bounds (2.3)-(2.5), that is, how much can we lose when we apply triangle inequality and Thm.
-1.2 in the proof of Thm.
-2.1?
-The loss can be significant: substitute $\rho=r$ into (2.3) and obtain $\mu:=||M-M(\rho)_{r}||\leq(1+\beta)\sigma_{r+1}(M)$, but more directly obtain $\mu:=||M-M(\rho)_{r}||=||M-M(r)||=\beta\sigma_{r+1}(M)$ because $M(\rho)_{r}=M(r)$.
-3 Iterative escalation with compression
-Next we elaborate upon iterative escalation with compression, sketched in Sec. 1.3.4.
+Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$, but can be very pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
+Remark 2.2.
+[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should keep $\rho$ small enough to perform at stage 2 fast or superfast.
+For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
+Remark 2.3.
+[A simplified version.] In Sec.
+5.3 we test a simplified version of Alg.
+2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$.
+In practice, the choice of integers $\rho$ and then $r$ should adaptive, towards minimizing the error norms of the resulting rank-$\rho$ and rank-$r$ approximations of $M$.
+3 LRA by means of random sampling with sparse multipliers
+The random sampling LRA algorithms of [11, 33] are close to optimal with a high probability but are not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT.
+In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations to superfast level by using sparse test matrices, e.g., Ultra-Sparse Rademacher test matrices [33, Sec.
+3.9], [3], which depend on a positive integer parameter $s$, such that a test matrices are sparse and support superfast LRA for smaller values $s$.
+It is a challenge to support our VLRA with these algorithms and test matrices, but in this paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix D and also supporting superfast VLRA.
+For other candidate sparse test matrices see [11, Remark 4.6], [7, Sec.
+5.1.8].
+According to [13], such acceleration tends to make the output LRAs somewhat less reliable, and indeed in our tests LRAs computed at stage 1 of Alg.
+2.1 tended to be cruder with abridged SRHT test matrices than with Gaussian random ones.
+Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
+4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
+Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec.
+1.3.
+We allow at most $k$ refinement steps for a fixed tolerance $k$.
+As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
 Algorithm 4.1.
-Iterative escalation.
+Iterative refinement of LRA of an $(r,\rho)$-matrix.
 Input and OUTPUT
@@
 2.1, except that the input of Alg.
-3.1 additionally includes a positive integer $k$ and an Alg.
-$\mathbb{S}$ that adds or subtracts a pair of black box $m\times n$ matrices at a computational cost $\tau(\pm)$.
+4.1 additionally includes a positive integer $k$, which bounds the number of recursive steps allowed, and an Alg.
+$\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
 Initialization:
 Write $i:=0$, $X_{0}:=0$, and $E_{0}:=M$ and fix an upper rank $\rho$ satisfying
-$2r\leq\rho\leq n\leq m,~{}~{}(\tau(\rho)+\tau(\rho,r)+\tau(r^{\prime})+2\tau(\pm))~{}k\leq\widehat{\tau}.$ (3.1)
+$2r\leq\rho\ll n.$ (4.1)
 Computations:
   1. 1.
-    Apply Alg. ${\bf\mathbb{A}}$ to the matrix $E_{i}$ and the upper rank $\rho$. [The algorithm outputs a rank-$\rho$ approximation of $E_{i}$ given by SVD of a matrix $E_{i}(\rho)$.]
+    Compute an approximation of $E_{i}$ by SVD of a matrix $E_{i}(\rho)$ of rank $\rho$. [This is step 1 of Alg. 2.1 applied to $E_{i}$ rather than $M$.]
   2. 2.
-    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$.
+    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$ by applying Alg. $\mathbb{S}$, performed with a higher precision. [This updates approximation $X_{i}$ to $M$.]
   3. 3.
-    Apply Alg. ${\bf\mathbb{B}}$ to this matrix and the fixed $r$ to compute $r$-top SVD of $M_{i}(\rho)$, that is, SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$.
+    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This compresses $M_{i}(\rho)$.]
   4. 4.
-    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if Alg. ${\bf\mathbb{C}}$ shows that $2\mu_{i}:=2\mu_{i}(M,\rho)\leq\widehat{\mu}$.
-    Otherwise output FAILURE if $i=k-1$.
+    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if it satisfies the tolerance bound, that is, if the $\mu_{i}:=\mu_{i}(M,\rho)\leq\mu$.
+    Otherwise output FAILURE if $i=k$.
   5. 5.
-    Otherwise compute $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1.
+    Otherwise, update the error matrix $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1 (to resume the iteration).
 Remark 4.1.
-[The computational cost bound, the choice of an upper rank, and a simplified version of Alg.
-3.1.] (i) (3.1) implies that the overall computational cost $\tau(\rho)+\tau(\rho,r)+\tau^{\prime}$ stays within tolerance $\widehat{\tau}$.
-(ii) As in Alg.
-2.1 we maximize the upper rank $\rho$ to decrease the output error norm, being only limited by a bound on the overall computational cost.
-(iii) In principle, we can allow an upper rank $\rho$ to vary as $i$ varies; then we can readily extend the analysis of the algorithm.
-(iv) In Sec.
-6.3 we test a simplified version of Alg.
-3.1 where we impose no upper bounds $\widehat{\tau}$ and $\widehat{\mu}$, fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ elsewhere.
-In the following extension of Thm. 2.1 we estimate the output error norm of Alg. 3.1.
+[The computational cost bound, the choice of an upper rank, and variations of Alg.
+4.1.] As in Alg.
+2.1 maximize the upper rank $\rho$ to decrease the output error norm, being only limited by the assumption that the overall computational cost is dominated at stage 1 (cf.
+(4.1)).
+We can allow an upper rank $\rho$ to vary as $i$ varies as long that it stays much lower than $n$ for all $i$.
+In Sec.
+5.3 we test a variant of Alg.
+4.1 where we fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
+In the following extension of Thm. 2.3 we estimate the output error norm of Alg. 4.1.
 Theorem 4.1.
-For the error matrices $E_{i+1}$ of Alg. 3.1, $\nu_{i}:=||E_{i}-E_{i}(\rho)||$, $i=0,1,\dots,k-1$, and an upper bound
-$\beta\geq\beta_{i}(M,\rho):=\frac{\nu_{i}}{\sigma_{\rho+1}(E_{i})}~{}{\rm for}~{}i=0,1,\dots,k-1$ (3.2)
-on the relative errors of Alg. $\mathbb{A}$ (unlike (1.3) we write $\beta$ rather than $\beta_{+}$), it holds that
-$\mu_{i+1}:=||E_{i+1}||\leq||M-M_{r}||+\nu_{i}=\sigma_{r+1}(M)+\beta\sigma_{\rho+1}(E_{i}).$
+For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,k-1.$ Then it holds that
+$\mu_{i+1}:=||E_{i+1}||\leq\sigma_{r+1}(M)+2\nu_{i}.$ (4.2)
 Proof.
-Recall from Alg.
-3.1 that $E_{i+1}=M-X_{i+1}$ for $X_{i+1}=(X_{i}+E_{i}(\rho))_{r}$.
-Hence $E_{i+1}=M-M_{r}+M_{r}-X_{i+1}$, and so $||E_{i+1}||\leq||M-M_{r}||+||M_{r}-X_{i+1}||=\sigma_{r+1}(M)+||M_{r}-X_{i+1}||$.
-It remains to prove that
-$||M_{r}-X_{i+1}||\leq\beta\sigma_{\rho+1}(E_{i}).$ (3.3)
-Since $X_{i+1}=(X_{i}+E_{i}(\rho))_{r}$, it follows that $M_{r}-X_{i+1}=M_{r}-(X_{i}+E_{i}(\rho))_{r}$.
-Combine this equation with Thm.
-1.2 and obtain that $||M_{r}-X_{i+1}||\leq||M-(X_{i}+E_{i}(\rho))||=||E_{i}-E_{i}(\rho)||$.
-Now bound (3.3) and hence the theorem follow from Eqn.
-(3.2).
-∎
+Recall from Alg. 4.1 that
+$E_{i+1}:=M-X_{i+1}~{}{\rm for}~{}X_{i+1}:=(X_{i}+E_{i}(\rho))_{r}.$
+Hence
+$||E_{i+1}||\leq||M-M_{i}(\rho)||+||M_{i}(\rho)-X_{i+1}||=\nu_{i}+\sigma_{r+1}(M_{i}(\rho)).$
+Thm. 2.2 implies that
+$\sigma_{r+1}(M_{i}(\rho))\leq\sigma_{r+1}(M)+\nu_{i}.$
+Combine the latter two bounds. ∎
+Remark 4.2.
+The upper bound (4.2) on the error norm is within $2\nu_{i}$ from the lower bound $\sigma_{r+1}(M)$, attained where $M_{i}(\rho)=M_{\rho}$.
+Next we specify how compression limits the accuracy of the output LRAs.
+Lemma 4.1.
+For any $m\times n$ matrix $X$ and $m\geq n$ write $\sigma_{j}(X):=\infty$ for $j<1$ and $\sigma_{j}(X):=0$ for $j>n$. Then
+[…]
+Proof.
+Recall that
+$E_{i}=M-X_{i}~{}{\rm where}~{}\operatorname{rank}(X_{i})\leq r~{}{\rm for~{}all}~{}i.$
+Furthermore,
+$\sigma_{j+1}(E_{i})=\min_{Y:~{}\operatorname{rank}(Y)\leq j}||E_{i}-Y||$
+by virtue of Thm. 2.1.
+Substitute $E_{i}-Y=M-(Y+X_{i})$ where $\operatorname{rank}(Y+X_{i})\leq j+r$ if $\operatorname{rank}(Y)\leq j$ and obtain
+$\sigma_{j+1}(E_{i})\geq\min_{Z:~{}\operatorname{rank}(Z)\leq j+r}||M-Z||=\sigma_{j+r+1}(M).$
+Similarly deduce that $\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$. ∎
 Corollary 4.1.
-Extend Cor. 2.1 by writing $\alpha_{i+1}$ and $E_{i+1}$ instead of $\alpha$ and $E$. Then
-$\alpha_{i+1}:=\alpha_{i+1}(M,r,\rho):=\frac{||E_{i+1}||}{\sigma_{r+1}(M)}-1\leq\beta\frac{\sigma_{\rho+1}(E_{i})}{\sigma_{r+1}(M)},~{}i=0,1,\dots,k-1.$ (3.4)
-Lemma 3.1.
-Recall that $\sigma_{j}(X)=0$ for $j>n$ and write $\sigma_{j}(X)=\infty$ for any $m\times n$ matrix $X$ and $m\geq n$.
-Then $\sigma_{j+r+1}(M)\leq\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$ for $E_{i}$ of Alg.
-3.1 and all $i$ and $j$.
-Proof.
-Recall that $\operatorname{rank}(M-X_{i})\leq r$ for all $i\geq 0$, $\sigma_{j+1}(E_{i})$ is an optimal rank-$j$ approximation of the matrix $E_{i}$ by virtue of Thm.
-1.1, and rank-$\rho$ approximation of $M$ with $Y$ does not exceed rank-$(\rho-r)$ approximation of $M$ with $Y+X$ if $\operatorname{rank}(X)=r$.
-Combine these observations.
-∎
-Combine bound (3.4) with the lemma for $j=\rho+1$ to obtain
-Corollary 3.2.
-The singular values $\sigma_{\rho+1}(E_{i})$ of Eqn. (3.4) satisfy
+The singular values $\sigma_{\rho+1}(E_{i})$ satisfy
 $\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,k-1.$ (4.3)
 Remark 4.3.
-[Error norm bounds.] For $E_{i+1}=E$ and $E_{i}=M$, bound (3.4) turns into bound (2.5) of Cor.
-2.1, and this is precisely what we have at the initial stage for $i=0$.
-For $i>0$ bound (3.4) can be weaker or stronger than that, but it cannot be stronger than (2.5) with $\rho$ replaced by $\rho+r$.
-In our tests for Algs.
-2.1 and 3.1 in Sec.
-6.3 for for matrices Shaw and $r=20$ Alg.
-3.1 for $\rho=2r$ superseded Alg.
-2.1 for $\rho=ir$, $i=3,4,5$.
-This implies that the upper bound of Cor.
-3.2 is overly pessimistic for worst case inputs, similarly to bound of Cor.
-2.1 (cf.
-Remark 2.2).
-4 LRA by means of random sampling with sparse multipliers
-Alg. 1.1 is near-optimal with a high probability (whp) but not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT [20], [54, Sec. 3.9].
-One can accelerate it to a superfast level by using Ultra-Sparse Rademacher test matrices [54, Sec.
-3.9]), [6], which depend on an integer parameter $s$, defining sparsity.
-They are sparse and support superfast LRAs for smaller values $s$.
-Likewise, in Appendix D we turn SRHT test matrices into abridged SRHT test matrices of a small length, which support superfast LRA; this recipe can be extended to various other recursively generated test matrices such as SRFT matrices and those of [20, Remark 4.6], [15, Sec.
-5.1.8].
-According to [22], such acceleration tends to make the output LRAs somewhat less reliable.
-Apparently for this reason, our tests for synthetic inputs from [55, Sec.
-7.3.1] failed with abridged SRHT test matrices of length 3, while the same tests succeeded with Gaussian test matrices.
-For the other inputs, however, our tests succeeded about equally well with test matrices of both kinds, even where the output LRAs of Alg.
-$\mathbb{A}$ were much cruder with abridged SRHT test matrices than with Gaussian random ones.
-[40, 41, 42, 43] studied primitive algorithms for CUR LRAs (cf.
-Appendix C.1) generated by pairs of test matrices made up of two random subsets of $k$ rows and $\ell$ columns of the identity matrix.
-These algorithms run at very low cost and output rather crude LRAs, and so [40, 41, 42, 43] proposed to strengthen them with escalation, by computing the sets of $k$ rows and $\ell$ columns as proper subsets of the sets of $p>k$ random rows and $q>\ell$ random columns of the identity matrix.
-[40, 41, 42, 43] call these algorithms cynical – representing a minor advance from the primitive ones.
-Not competitive for LRA, such cynical algorithms were predecessors of our use of escalation and could still be considered for Alg.
-$\mathbb{A}$.
-5 Superfast 1-norm estimator
-Next we propose a superfast variation of the LAPACK estimator for 1-norm $||E||_{1}$ of a matrix $E\in\mathbb{R}^{m\times n}$ (cf.
-[19, Alg.
-15.3], [51, Alg.
-5.3.1]) and recall that $||E||_{1}=||E^{T}||_{\infty}$, $||E||_{\infty}\leq n||E||$, $||E||_{1}\leq m||E||$, and $||E||\leq||E||_{1}||E||_{\infty}$ (see [15, Eqns.
-(2.3.11), (2.3.12) and Cor.
-2.3.2])).
-Empirically, the LAPACKl norm-1 estimator converges in about 4.5 iterations on the average and is superfast at every iteration, except for its initial and verification stages, at which $E$ is multiplied by dense auxiliary vectors.
-We accelerate these stages by using sparse vectors instead.
-We keep a small number of randomly sampled components of the original vectors, set the remaining components to 0, and scale the resulting vectors to turn their 1-norm into 1; we only test our 1-norm estimator for matrices $E$ representing the errors of LRA.
-Algorithm 5.1.
-[A superfast 1-Norm Estimator.]
-Input:
-An $m\times n$ matrix $E$, a positive integer $k\ll n\leq m$, and a positive tolerance TOL.
-Output:
-A vector ${\bf u}$ such that $||E{\bf u}||_{1}\approx||E||_{1}$.
-Initialization:
-Generate a sparse unit vector ${\bf v}$ such that $||{\bf v}||_{1}=1$ as follows: choose $k$ random integers $i_{1},\dots,i_{k}$ in the range from 1 to $n$ and let ${\bf v}=\frac{1}{k}(v_{i})_{i=1}^{n}$ where $v_{i_{j}}=1$ for $j=1,\dots,k$ and $v_{i}=0$ for all other $i$.
-Write COUNTER:=0.
-Computations:
-  1. 1.
-    Compute the vectors ${\bf u}=E{\bf v}$, ${\bf w}=(\operatorname{sign}(u_{i}))_{i=1}^{n}$, and ${\bf x}=E^{T}{\bf w}$.
-  2. 2.
-    If $||{\bf x}||_{\infty}\leq||{\bf u}||_{1}$, output ${\bf u}$ and terminate.
-  3. 3.
-    Choose $j$ such that $|x_{j}|=||{\bf x}||_{\infty}$ and let ${\bf v}={\bf e}_{j}$. Increase COUNTER by 1. Go to stage 1 if COUNTER$<$Tol. Otherwise stop and output FAILURE.
-We also apply the algorithm to a small number of other choices of the initial sparse unit vector ${\bf v}$ and output ${\bf u}$ that maximizes the norm $||E{\bf u}||_{1}$.
-For these choices we first define vectors of dimension $n$ sharing $k$ components on $k$ random positions with the following vector
-$\left(1,-1-\frac{1}{n-1},1+\frac{2}{n-1},1-\frac{3}{n-1},\dots,(-1)^{n-1}2\right)^{T}$
-and otherwise filled with 0s; then scale them to make their 1-norm equal to 1.
+[Error norm bounds.] In good accordance with the corollary, in our tests with $(r,\rho)$-matrices the output accuracy of Alg.
+2.1 was higher than that of Alg.
+4.1 provided that the upper rank $\rho$ in Alg.
+2.1 was twice as large as in Alg.
+4.1.
+The test results were mixed for input matrices with flat spectra.
 5 Numerical experiments
-We devote the next three subsections to the tests of our LRA algorithms on synthetic and real-world input matrices devote Sec. 6.4 to testing Alg. 5.1.
-We implemented our algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory.
-We called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
+In this section we cover the tests of our LRA algorithms on synthetic and real-world matrices.
+We implemented the algorithms in Python, running them on a 64bit MacOS machine with 16GB Memory and called scipy.linalg for numerical linear algebra routines such as QR factorization with pivoting, Moore-Penrose matrix inversion, and SVD.
 5.1 Input Matrices
 […]
-Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [20, Sec. 7.1] for more details).
-We display the distribution of the top 50 singular values of these three matrices in Fig. 1.
+Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [11, Sec. 7.1] for more details).
+Fig. 1 shows the distribution of the top 50 singular values of these three matrices.
 We set the target ranks $r$ equal to 20 for Shaw, 45 for Gravity, and 11 for SLP matrices.
-Figure 1: Spectra of singular values of some real world matrices
-Furthermore, we computed rank-$r$ approximations for the three matrices, StreamVel, MinTemp, and MaxCut, used in numerical experiments in [55, Sec. 7.3.2].
+Figure 1: Spectra of singular values of some real-world matrices
+We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [34, Sec. 7.3.2].
 The $10738\times 5001$ matrix StreamVel is involved in numerical simulation (DNS) on a coarse mesh of the 2D Navier–Stokes equations for a low-Reynolds number flow around a cylinder.
@@
 The $2000\times 2000$ matrix MaxCut provides a high-accuracy solution to the MaxCut SDP for a sparse graph [9].
-In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2 and set the target rank $r$ equal to 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
+We let the target rank $r$ be 20 for StreamVel, 10 for MinTemp, and 14 for MaxCut matrices.
+In our tests, we padded all matrices with 0s to increase their dimensions to the powers of 2.
 Synthetic Input Matrices.
@@
 By letting $\Sigma=\textrm{diag}(u)$, $u_{i}=1$ for $i=1,2,\dots,20$, and $u_{i}=\frac{1}{(1+i-20)^{2}}$ for $i>20$, we arrived at the matrices with slowly decaying spectra.
-For both input classes we fixed target rank $r=20$ and set it equal to 10 for the next three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [55, Sec.
+For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [34, Sec.
 7.3.1], which involved effective rank parameter $R=20$.
-(i) Low-rank + noise matrices
-${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$
-for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, also involves parameter $\xi$, equal to $10^{-4}$ for the Low Rank Low Noise matrix, to $10^{-2}$ for the Low Rank Med Noise matrix, and to $10^{-1}$ for the Low Rank High Noise matrix.
-(ii) We generated Polynomial Decay matrices
-${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p}),$
-where $p=0.5$ for Poly Decay Slow, $p=1$ for Poly Decay Med, $p=2$ for Poly Decay Fast.
-(iii) We generated Exponential Decay matrices
-${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q}),$
-with $q=0.01$ for Exp Decay Slow, $q=0.1$ for Exp Decay Med, $q=0.5$ for Exp Decay Fast.
-6.2 Two options for Alg. $\mathbb{A}$
-In our tests of Algs. 2.1 and 3.1 we used the following two options for Alg. $\mathbb{A}$, which computes a crude rank-$\rho$ approximation of a matrix $M$.
-(i) Fast (although not superfast) randomized algorithm of [54] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times(\rho)$.
-(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{R}^{(2\rho)\times 1024}$ and $H\in\mathbb{R}^{1024\times\rho}$ replaced Gaussian ones.
-6.3 Test results for Algs. 2.1 and 3.1
-In our tests of escalation Alg. 2.1 we output the relative error norms
-$\text{err}=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$
+(i) Low-rank + noise matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},0,\dots,0)+(\xi/1024){\bf C},$ for ${\bf C}={\bf G}{\bf G}^{T}$ and a $1024\times 1024$ standard normal matrix ${\bf G}$, depend on a parameter $\xi$ and become Low Rank Low Noise matrix for $\xi=10^{-4}$, the Low Rank Med Noise matrix for $\xi=10^{-2}$, and the Low Rank High Noise matrix for $\xi=10^{-1}$.
+(ii) Polynomial Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},2^{-p},3^{-p},\dots,(n-R+1)^{-p})$ are Poly Decay
+Slow
+for $p=0.5$, Poly Decay Med for $p=1$, and Poly Decay Fast for $p=2$.
+(iii) Exponential Decay matrices ${\bf A}=\operatorname{diag}(\underbrace{1,\dots,1}_{R},10^{-q},10^{-2q},\dots,10^{-(n-R)q})$ are Exp Decay
+Slow
+for $q=0.01$, Exp Decay Med for $q=0.1$, and Exp Decay Fast for $q=0.5$.
+5.2 Two options at stages 1 of Algs. 2.1 and 4.1
+In our tests of Algs. 2.1 and 4.1 we used the following two options at their stage 1 for computations of a crude rank-$\rho$ approximation of a matrix $M$.
+(i) Fast (although not superfast) randomized algorithm of [33] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
+(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{C}^{(2\rho)\times 1024}$ and $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
+5.3 Test results for Algs. 2.1 and 4.1
+For Alg.
+2.1 we output the relative error norms $\text{err}:=\frac{||M-(M(\rho))_{r}||}{||M-M_{r}||}$ for input matrices $M$, a target rank $r$, the $r$-truncation $M_{r}$ and $M(\rho)_{r}$, and the upper rank $\rho$, set to $2r,3r,4r$, $5r$ (cf.
+Remark 2.3).
 […]
-In our tests of iterative escalation Alg. 3.1 we output the relative error norms
-[…]
-[…]
-We only display the error norms output in the first three iterations of Alg.
-3.1 and notice that, except for the matrices representing high noise or slow decay, the second iteration greatly decreased the output error norms, while the third iteration tended to keep the error norms at more or less the same level.
-[…]
-For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple of $r$, $\rho$, and an input matrix type for Alg.
-2.1, and per the pair of $r$ and an input matrix type for Alg.
-3.1, except that we ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger size.
 Our tables display the mean of the relative error norms observed in these tests.
 For Alg. 2.1 we also display the standard deviation.
-In Tables 6.1, 6.2, 6.5, and 6.6 we display the results of our tests with both Gaussian random and abridged SRHT test matrices.
-This covers the cases where the output error norms were more or less in the same range for both test matrices (and usually quite similar in both cases, with only a slightly better results with Gaussian random test matrices).
-Displaying the results of our tests with the synthetic input matrices from [55, Sec.
-7.3.1] in Tables 6.3 and 6.4, however, we only show the test results with Gaussian random test matrices because all tests with abridged SRHT test matrices have failed, that is, have output with large errors.
-The tests with Gaussian random test matrices have output accurate LRAs for the same input matrices, except for those representing high noise or slow decay.
+For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg.
+2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg.
+4.1 where TYPE stands for an input matrix type.
+We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
+In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT test matrices, which were more or less in the same range for both matrix types.
+All our tests with the synthetic input matrices from [34, Sec.
+7.3.1] wit abridged SRHT test matrices have failed, that is, have output large error norms.
+In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random test matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
 Alg.
@@
 The relative error norms output by both Algs.
-2.1 and 3.1 have consistently decreased and frequently to the optimal value 1 (up to within the measurement errors) as the spectra of the input matrices decayed.
-The error norms have also decreased as we increased the upper rank $\rho$ for Alg.
-2.1 and, as we said already, in the transition from the first to the second iteration of Alg.
-3.1.
-The error norm bounds of Alg.
+2.1 and 4.1 applied to $(r,\rho)$-matrices have consistently decreased and frequently to the optimal value 1 (up to the measurement errors) as we increased the upper rank $\rho$ for Alg.
+2.1 and in the transition from the first to the second iteration of Alg.
+4.1.
+In all tests except for High Noise, Slow Decay, and MinTemp matrices already err2 decreased below 1.1.
+The third iteration decreased the error norm much less or even slightly increased it.
+For $(r,\rho)$-matrices the error norm bounds of Alg.
 2.1 for $\rho=4r$ were consistently smaller than those of Alg.
-3.1 in all three iterations, except for tests with the matrix Shaw for $r=20$.
-This exception implies that our estimate of Cor.
-3.2 is overly pessimistic for worst case inputs, as we explained in Remark 3.2.
+4.1 in all three iterations.
+This was not the case just for the matrices with flat spectra such as the matrices Shaw for $r=20$, High Noise, and Exp Decay Slow.
 Input Matrix Multiplier $\rho$
@@
 Gaussian 1.000 $\pm$ 4.130E-05 1.000 $\pm$ 1.380E-06 1.000 $\pm$ 1.783E-07 1.000 $\pm$ 3.708E-08
-Table 6.1: Test results for Alg.
-2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, synthetic matrices with fast decaying spectra (Fast Decay), and synthetic matrices with slow decaying spectra (Slow Decay).
+Table 5.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, and synthetic matrices Fast Decay and Slow Decay, with fast and slowly decaying spectra.
 Input Matrix Multiplier 1st Itr. 2nd Itr. 3rd Itr.
@@
 Gaussian 5.2205 1.0000 1.0000
-Table 6.2: Test results for Alg.
-3.1 for $\rho=2r$ on synthetic matrices with fast decaying spectra (Fast Decay), synthetic matrices with slow decaying spectra (Slow Decay), Shaw, Gravity, and discretized single layer potential operator (SLP).
+Table 5.2: Test results for Alg. 4.1 on Fast Decay, Slow Decay, Shaw, Gravity, and SLP matrices.
 Input Matrix $\rho$
@@
 Exp Decay Fast 1.3121 $\pm$ 1.4989E-01 1.$0$$\pm$ 6.4663E-11 1.$0$$\pm$ 3.6020E-16 1.$0$$\pm$ 3.0986E-16
-Table 6.3: Test results for Alg. 2.1 on the synthetic input matrices from [55, Sec. 7.3.1] using Gaussian multipliers.
+Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian test matrices.
 Input Matrix 1st Itr. 2nd Itr. 3rd Itr.
@@
 Exp Decay Fast 1.4235 1.0000 1.0000
-Table 6.4: Test results for Alg. 3.1 on the synthetic input matrices from [55, Sec. 7.3.1] using Gaussian multipliers.
+Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian multipliers.
 Input Multiplier $\rho$
@@
 Gaussian 1.0026 $\pm$ 1.3856E-03 1.0005 $\pm$ 2.1129E-04 1.0002 $\pm$ 6.6260E-05 1.0001 $\pm$ 2.4207E-05
-Table 6.5: Test results for Alg. 2.1 on the real world input matrices from [55, Sec. 7.3.2]
+Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [34, Sec. 7.3.2]
 Input Matrix Multiplier Type 1st Itr. 2nd Itr. 3rd Itr.
@@
 Gaussian 2.6202E+01 1.0302 1.0176
-Table 6.6: Test results for Alg. 3.1 for $\rho=2r$ on real world input matrices from [55, Sec. 7.3.2].
-6.4 Test results for Alg. 5.1
-For each $n\times n$ matrix $M$, we computed its rank-10 approximation $\tilde{M}$ by its truncated SVD and then obtained the error matrix $E=M-\tilde{M}$.
-We generated initial vectors ${\bf v}$ according to the following
-INITIALIZATION RULE: Define a vector ${\bf v}$ of length $n$ with coordinates $1/k$ at $k$ positions sampled at random independently under the uniform probability distribution; fill the vector with 0s elsewhere; do this for $k=1,\log\log(n),\log(n)\text{ and }n$.
-We performed 100 tests for each pair of input matrix type and integer $k$ and measured the accuracy of the estimates by the ratio
-$\text{accuracy}=\frac{||E||_{1}}{\text{est}},$
-where est denoted the estimate of $||E||_{1}$ returned by the algorithm.
-The ratio took on its minimal value 1 where the estimate est was optimal.
-In our test most of the norm estimates est returned by the algorithm were within a factor of 2 from the optimal norm $||E||_{1}$, even where the initial vector had just a single nonzero entry.
-All our tests converged within 6 iterations, regardless of the density of the initial vector.
+Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [34, Sec. 7.3.2].
 6 Conclusions
-We proposed, tested, and analyzed some simple but novel LRA algorithms using escalation and proved that at a dominated or even sublinear additional computational cost they significantly decrease the spectral norms of the output error matrices of the known algorithms applied to FDS matrices.
-The results of our numerical tests with a variety of synthetic and real world matrices are in good accordance with our formal estimates.
-Furthermore, some superfast (sublinear cost) variants of the known algorithms are known to output LRA that tend to be less accurate.
-In combination with our algorithms, however, this deficiency has been fixed in most of our tests, and it is a challenge to test the output accuracy of these and other superfast (sublinear cost) LRA algorithms for a greater variety of inputs.
-Towards enhancing our progress to a larger class of matrices $M$, we nontrivially extended to LRA the popular technique of iterative refinement.
-So far the resulting algorithm was superior to our escalation LRA for only a single input class of matrices among 17 classes tested, but further comparative study may still be in order.
-Another natural but yet unexplored challenge is to compare and possibly combine the iterative refinement LRA algorithm of [25] with our present escalation and iterative refinement algorithms.
+Rank-$r$ approximation of an $m\times n$ matrix $M$ is only interesting if there is a gap in the spectrum of its top (largest) $r+1$ singular values $\sigma_{j}(M)$.
+For the $(r,\rho)$-matrices $M$ having also a gap in the next segment of their spectra – from $(r+1)$st to $(\rho+1)$st largest singular values – randomized oversampling algorithms output quite accurate LRA.
+We extended these observations towards devising superfast (sublinear cost) variation of LRA algorithms based on random sampling and oversampling with sparse test matrices and to improving accuracy of the resulting LRA in the case of $(r,\rho)$-matrices $M$ where $r<\rho\ll\sqrt{n}$.
+We provided some formal and experimental support for the efficiency of the resulting LRA algorithms.
+There are various natural extensions of our study, which can be practically competitive.
+E.g., one can devise and test similar variations of the TYUC and GN algorithms with sparse test matrices such as Ultra-Sparse Rademacher test matrices [33, Sec.
+3.9] or sparsified random Givens rotations by Rokhlin and Tygert (see [11, Remark 4.6]).666[26, 27, 28] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes.
+Our study in Secs. 4 and 5 has showed no benefits for using iterative refinement of LRA versus oversampling so far, but one can keep comparing their efficiency.
 Appendix
@@
 Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank.
-Furthermore, the same argument shows that superfast estimation of the norm of a matrix fails already on the same matrix family and that for both problems of LRA and norm estimation randomized superfast algorithms fail with error probability not close to 0.
+Likewise, we can verify that any superfast randomized LRA algorithms fail fails on those matrix families with error probability not close to 0.
 Remark A.1.
@@
 Appendix B Computation of $r$-top SVD of a matrix product
-For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [31, 32, 37, 46, 38]) based on computation of the SVDs of $A$ and $B$.
+For completeness of our exposition we next recall the classical algorithm that computes $r$-top SVD of a matrix product $AB$ (cf. [22, 30]) based on computation of the SVDs of $A$ and $B$.
 Algorithm B.1.
@@
 Output:
-Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{R}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
+Three matrices $U\in\mathbb{C}^{m\times r}$ (unitary), $\Sigma\in\mathbb{C}^{r\times r}$ (diagonal), and $V\in\mathbb{C}^{n\times r}$ (unitary) such that $(AB)_{r}=U\Sigma V^{*}$ is an $r$-top SVD of $AB$.
 Computations:
@@
 The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf.
-[15, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$.
+[7, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$.
 Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$.
@@
 $R:=M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},C:=M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{}{\rm and}~{}G:=M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}$
-of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$ (cf.
-[8]), although we recall notation $G$ for nucleus and CGR for the factorization in [30], originated with the decomposition itself in the 1990s [17].
+of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$, although we recall notation $G$ for nucleus and CGR for the factorization in [21], originated with the decomposition itself in the 1990s [8].
 In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
@@
 (See [28].) For matrices $C$, $U$, and $R$ defined above, $M=CUR$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
-If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large, but empirically the Cross-Approximation (C-A) celebrated iterations, briefly recalled in Appendix E, tend to converge fast to quite an accurate rank-$r$ CUR approximation for a large class of matrices $M$.
+If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large.
 C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
@@
 Input:
-Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{R}^{m\times n}$, $\Sigma\in\mathbb{R}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{R}^{m\times r}$, and $V\in\mathbb{R}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
+Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{C}^{m\times n}$, $\Sigma\in\mathbb{C}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{C}^{m\times r}$, and $V\in\mathbb{C}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
 Output:
-Three matrices $C\in\mathbb{R}^{m\times r}$, $N\in\mathbb{R}^{r\times r}$, and $R\in\mathbb{R}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
+Three matrices $C\in\mathbb{C}^{m\times r}$, $N\in\mathbb{C}^{r\times r}$, and $R\in\mathbb{C}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
 $M=CNR.$
@@
 Appendix D Generation of abridged SRHT matrices
-In this section we specify the family of abridged SRHT matrices, used in Alg.
-1.1.
+In this section we specify the family of abridged SRHT matrices, used in our tests for Algs.
+2.1 and 4.1.
 They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf.
-[26, Sec.
+[17, Sec.
 3.1]).
@@
 The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf.
-[20, Sec.
+[11, Sec.
 11]).
-In this Appendix we follow [40, 41, 43, 42] where abridged SRFT matrices have been defined and tested similarly to abridged SRHT matrices with similar test results.
-Appendix E Cross-Approximation (C-A) iterations
-In this appendix we briefly recall C-A iterations, which can be viewed as a specialization to LRA of Alternative Directions Implicit (ADI) method.
-C-A iterations:
-  • For an $m\times n$ matrix $M$ and a target rank $r$, fix four integers $k$, $\ell$, $p$ and $q$ such that $0<r\leq k\leq p\leq m$ and $r\leq\ell\leq q\leq n$.
-[C-A iterations are simplified if $p:=k$ and $q:=\ell$.]
-  • Fix an $m\times q$ “vertical” submatrix of the matrix $M$, made up of its $q$ columns. [^14] Of course, one can alternatively begin C-A iterations with a “horizontal” submatrix.
-  • By applying a fixed CUR LRA subalgorithm, e.g., one of the algorithms of [29, 14, 33, 8], [^15] Such a subalgorithm runs at sublinear cost on the inputs of smaller size involved in C-A iterations, even where the original algorithms of [29, 14, 33, 8] applied to an $m\times n$ matrix $M$ run at superlinear cost.
-compute a $k\times\ell$ CUR generator $G$ of this submatrix and reuse it for the matrix $M$.
-  • Compute the rank-$r$ CUR of $M$ based on the generator $G$ and output this CUR it if it is close enough.
-  • Otherwise swap $p$ and $q$ and reapply the algorithm to the matrix $M^{*}$.
-    [This is equivalent to computing a $k\times\ell$ CUR generator of a fixed $p\times n$ “horizontal” submatrix $M_{1}$ of $M$ that covers the submatrix $G$.]
-  • Recursively alternate such “vertical” and “horizontal” steps until an accurate CUR LRA is computed or until the number of recursive C-A steps exceeds a fixed tolerance bound.
-Figure 2: The first three recursive C-A steps output three striped submatrices.
-Initialization recipes.
-For a large class of input matrices it is efficient to apply adaptive C-A iterations (cf.
-[1, 2]).
-They combine initialization with dynamic search for gaps in the singular values of $M$ by properly adapting Gaussian elimination with pivoting.
-The alternative initialization in [16] uses $O(nr^{2})$ flops to initialize the C-A iterations for an $n\times n$ input and $q=s=k=\ell=r$, and then the algorithm uses $O(rn)$ flops per C-A step.
-The output of C-A iterations, however, can be off by a significant factor.
-Appendix F Iterative refinement of LRA with no compression
-The definition of $r$-truncation immediately implies
-Lemma F.1.
-$M_{s}+(M-M_{s})_{r}=M_{r+s}$ for any pair of positive integers $r$ and $s$.
-Next, based on this lemma, we lift $r$-top SVD into $kr$-top SVD by applying iterative refinement of LRA with no compression.
-Algorithm F.1.
-[Lifting $r$-top SVD into $kr$-top SVD.]
-Input:
-An $m\times n$ matrix $M$, $n\leq m$, two positive integers $k$ and $r$, an Alg.
-$\mathbb{S}$ that adds or subtracts a pair of $m\times n$ matrices, and an Alg.
-$\mathbb{T}$ that computes the $r$-truncation of an $m\times n$ matrix.
-Output:
-The $kr$-truncation of $M$.
-Initialization:
-Write $X_{0}:=0$ and $E_{0}:=M$.
-Computations:
-Recursively combine Algs. $\mathbb{S}$ and $\mathbb{T}$ to compute the matrices
-$X_{i+1}:=X_{i}+(E_{i})_{r},~{}E_{i+1}:=M-X_{i+1},~{}~{}i=0,1,\dots,k-1.$ (F.1)
-Output the matrix $X_{k}$.
-The following proposition implies correctness of the algorithm.
-Proposition F.1.
-It holds that $X_{i}=M_{ir}$ for the matrices M and $X_{i}$ of Alg. F.1 for $i=0,1,\dots,k$.
-Proof.
-The claim holds for $i=0$ because $M_{0}=0$.
-Let it hold for $i\leq j$, that is, let $X_{j}=M_{jr}$, and then deduce from Eqn.
-(F.1) for $i=j$ that $X_{j+1}=X_{j}+(M-X_{j})_{r}$.
-By assumption, $X_{j}=M_{jr}$.
-Combine the latter equations and obtain $X_{j+1}=M_{jr}+(M-M_{jr})_{r}$.
-Apply Lemma F.1 for $s=jr$ and obtain that $X_{j+1}=M_{(j+1)r}$.
-Now the proposition follows by induction.
-∎
-Trying to extend our error analysis of Alg.
-3.1 to Alg.
-F.1 we needed to estimate for the norm of the perturbation of the $r$-top SVD, but at this point the known studies (cf.
-[15, Thm.
-8.6.5], [52, Thm.
-3.3.4]) provided much weaker results than Thm.
-1.2, and we ended with no support for numerical stability of Alg.
-F.1.
+Acknowledgement: Jianlin Xia helped place our study into proper historical context.

### v17

@@
-Near-Optimal Very Low Rank Approximation at Sublinear Cost
+Iterative Refinement and Oversampling for Low Rank Approximation
 Abstract
-Eckart-Young’s theorem implies that even optimal rank-$r$ approximation of an $m\times n$ matrix $M$ for $m\geq n$ is only interesting if there is a gap in the segment of the $r+1$ largest singular values of $M$.
-If $M$ also has a gap in the segment of the next $p$ largest singular values, then one can compute its rank-$r$ approximation with random sampling and oversampling in nearly linear arithmetic time.
-For matrices $M$ having such a double gap in their spectra and for $n$ greatly exceeding $(r+p)^{2}$, our algorithm runs in sublinear time, and for a large subclass of the class of such matrices $M$ its output rank-$r$ approximation of $M$ is near-optimal under the spectral norm.
-We also extend the classical iterative refinement to LRA and analyze and test the resulting algorithm.
+Iterative refinement is particularly popular for numerical solution of linear systems of equations.
+We extend it to Low Rank Approximation of a matrix (LRA) and observe close link of the resulting algorithm to oversampling techniques, commonly used in randomized LRA algorithms.
+We elaborate upon this link and revisit oversampling and some efficient randomized LRA algorithms.
+Applied with sparse sketch matrices they run significantly faster and in particular yield Very Low Rank Approximation (VLRA) at sublinear cost, using much fewer scalars and flops than the input matrix has entries.
+This is achieved at the price of deterioration of output accuracy, but according to our formal and empirical study subsequent oversampling improves accuracy to near-optimal level under the spectral norm for a large sub-class of matrices with fast decaying spectra of singular values.
 Key Words:
-low rank approximation, singular values, gaps in the spectrum, sublinear cost algorithms, iterative refinement
+low rank approximation, iterative refinement, oversampling, sparse sketches, singular values, gaps in the spectrum, sublinear cost.
 2020 Math. Subject Classification:
 65F55, 65N75, 65Y20, 68Q25, 68W20
-1 LRA of a matrix: the State of the Art and our Novelties
-Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra.
-We first briefly recall the State of the Art, referring to [11, 33, 18] for broader expositions, and then outline our LRA algorithms.
-1.1 The State of the Art
+1 Introduction: LRA of a matrix, its iterative refinement, link to oversampling, and sparse sketches
+1.1 Brief overview
+Classical iterative refinement (see pointers to the bibliography in Sec.
+1.6) is particularly popular for numerical solution of linear systems of equations.
+We naturally extend it to Low Rank Approximation of a matrix (LRA), observe its close link to LRA with oversampling, and analyze both refinement and oversampling.
+Here is a brief outline of our study.
+Let an LRA algorithm $\mathbb{A}=\mathbb{A}_{r}$ be applied to an $m\times n$ matrix $E_{0}:=M$ and let it output a rank-$r$ approximation $Y_{0}:=X_{0}:=\mathbb{A}(E_{0})$ of $M$.
+[^1] Here and hereafter “rank-$r$ approximation” means “rank at most $r$ approximation”.
+Apply the algorithm $\mathbb{A}$ to the error matrix $E_{1}=E_{0}-X_{0}$ and add $X_{1}:=\mathbb{A}(E_{1})$ to $X_{0}$ to obtain improved approximation $Y_{1}:=X_{0}+X_{1}$ to $M$.
+By repeating this step recursively $h$ times, compute the matrices $X_{i}:=\mathbb{A}(E_{i})$, $Y_{i}:=\mathbb{A}(Y_{i-1}+X_{i})$, and $E_{i+1}:=E_{i}-X_{i}$, for $i=1,2,\dots,h$.
+Finally output the matrix $\mathbb{A}(Y_{h})$ as a rank-$r$ approximation of $M$.
+The compression of higher rank approximation of a matrix into its lower rank approximation also occurs in the customary random oversampling for LRA.
+We revisit random sampling LRA algorithms with oversampling, accelerate them by using sparse sketch matrices and in particular yield Very Low Rank Approximation (VLRA) at sublinear cost.
+We achieve this at the price of some deterioration of the accuracy of the output LRA, but according to our formal and empirical study, subsequent oversampling improves accuracy to near-optimal under the spectral norm for a large sub-class of the matrices with fast decaying spectra of singular values.
+So far, our study has showed no additional benefits for LRA from application of iterative refinement versus standalone oversampling.
+1.2 The State of the Art of LRA
+LRA is a hot topic of Numerical Linear Algebra. We first briefly recall the State of the Art, referring to [10, 27, 15] for broader expositions, and then outline our results for LRA.
 1.1.1.
-LRA and CUR LRA.
+Representation and the power of LRA.
 Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors.
-[^1] The sketch of a matrix in [36, 33] is its product with a fixed set of vectors.
-Then we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [33, Sec.
+[^2] The sketch of a matrix in [30, 27] is its product with a matrix of smaller size, that is, with a small set of vectors.
+In this way we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [27, Sec.
 1.3]).
-Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [35], that is, admit LRA, with which one can operate superfast, aka at sublinear cost – by using much fewer than $mn$ scalars and flops.
-[^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [24], and we extend this concept to other matrices.
-Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r\ll n$ or a fixed target range [^3] Here and hereafter $a\ll b$ means that $b$ greatly exceeds $a$ in context, e.g., $a=o(b)$ as $b\mapsto\infty$.
-for such a rank (actually, a proper choice of such targets is a challenge [11, Sec.
-4.2]) and then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
+Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [29], that is, admit LRA, with which we can operate efficiently, even where $M$ has enormous size.
+Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r<n$ or a fixed target range for such a rank, although actually a proper choice of such targets is a challenge (see Remark 1.1).
+Then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
 $X\in\mathbb{C}^{m\times k},Y\in\mathbb{C}^{k\times\ell},~{}Z\in\mathbb{C}^{\ell\times n},$ (1.1)
-$k\ll m,~{}\ell\ll n,~{}\min\{k,\ell\}=r,$ (1.2)
+$k\leq m,~{}\ell\leq n,~{}\min\{k,\ell\}=r,$ (1.2)
 and the spectral and/or Frobenius norm of the matrix $M-XYZ$ is small in context, say, close to optimal.
-The matrices $X,Y,Z$ have $mk+k\ell+\ell n\ll mn$ entries overall, and we can multiply the matrix $XYZ$ by a vector by using less than $2(mk+k\ell+\ell n)\ll mn$ flops.
-If the matrices $X$ and $Z$ are made up of the columns and rows of $M$, that is, are defined by means of pre- and post-multiplication of $M$ by selected coordinate vectors, respectively, then write $X:=C$, $Z:=R$, and $Y:=U$ and call the triplet $\{X,Y,Z\}$ a CUR LRA.
 1.1.2.
-LRA from top SVD, SVD of an LRA, and CUR from SVD.
-By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, obtain a matrix $M_{r}$ of rank $r$, said to be $r$-truncation of $M$.
-By virtue of the Eckart-Young-Mirsky theorem [7, Thm.
-2.4.8], $M_{r}$ is an optimal rank-$r$ approximation444“Rank- $r$ approximation” means “rank at most $r$ approximation” throughout.
-of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
-Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [7, Fig.
+LRA from top SVD; SVD of an LRA.
+By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, turn SVD of $M$ into its $r$-top SVD and turn $M$ into its $r$-truncation $M_{r}$, of rank $r$.
+By virtue of the Eckart-Young-Mirsky theorem [6, Thm.
+2.4.8], $r$-top SVD defines an optimal rank-$r$ approximation of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
+Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [6, Fig.
 8.6.1]).
-This is too expensive for computing LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ and then its CUR LRA by using $(m+k)k+(n+\ell)\ell$ scalars, $O(mk^{2}+n\ell^{2})$ flops and $(m+n+r)r$ scalars, $O(mr^{2})$ flops, respectively (see Appendices B and C.2, respectively).
+This is too expensive for LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ by using just $(m+k)k+(n+\ell)\ell$ scalars and $O(mk^{2}+n\ell^{2})$ flops (see Appendix B).
 1.1.3.
 LRA via random sampling and oversampling.
-Random sampling with oversampling enables fast computation of LRA (see [3, 17, 11, 36, 33, 20], and the references therein).
-For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [11] successively generates a random test matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
+Random sampling with oversampling enables fast computation of close LRA (see [2, 14, 10, 30, 27, 16], and the references therein).
+For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [10] successively generates a random sketch matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
 Remark 1.1.
 “In practice, the target rank $k$ is rarely known in advance.
-Randomized algorithms are usually implemented in anadaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [11, Sec.
+Randomized algorithms are usually implemented in an adaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [10, Sec.
 4.2]).
-In the best studied case the test matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables.
-In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [11, Thm.
+In the best studied case the sketch matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables.
+In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [10, Thm.
 10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops.
-The parameter $p$ and the size of the test matrix $H$ explode as $\epsilon\mapsto 0$, but LRA with $\epsilon$ as large as 1/2 or 1 is still valuable.
-In other LRA algorithms (e.g., in [4, 3, 33, 20]) $M$ is both pre- and post-multiplied by random test matrices $F$ and $H$, respectively.
-In particular, the TYUC algorithm of [33], modifying those of [3, Thm.
-4.7] and [3, Thm.
-4.3, display 1], and the generalized Nyström approximation GN of [20] first compute the products $FM$ and $MH$ for random test matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$ with the expected Frobenius error norm within a factor of 2 from optimal in the case of random Gaussian test matrices $F$ and $H$ (namely, the norm does not exceed the optimal one for rank-$\frac{k}{2}$ approximation).
+The integer $p$ and the size of the matrix $H$ explode as $\epsilon\mapsto 0$, but LRA for $\epsilon$ as large as 1/2 or 1 is still valuable.
+In other LRA algorithms (e.g., in [4, 2, 27, 16]) $M$ is both pre- and post-multiplied by random sketch matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$, respectively.
+In particular, the TYUC algorithm of [27], modifying those of [2, Thm.
+4.7] and [2, Thm.
+4.3, display 1], and the generalized Nyström approximation GN of [16] first compute the products $FM$ and $MH$ for random sketch matrices $F$ and $H$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$.
+In the case of random Gaussian sketch matrices $F$ and $H$ the expected Frobenius error norm of the output LRA is within a factor of 2 from optimal (namely, it does not exceed the optimal one for rank-$\frac{k}{2}$ approximation).
 For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [16]) and $O(\ell mn)$ flops.
-The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [11, Secs. 10 and 11], [36, Ch. 6.2], [5, Thm. 3.4]).
-1.1.4.
-Superfast LRA for special subclasses of matrices.
-For some special but large and important classes of matrices, most notably for Symmetric Positive Semidefinite (SPSD) matrices but also for Distance matrices and a subclass of Cauchy matrices, the superfast randomized algorithms of [19, 1, 12, 14, 2] compute LRA whose expected Frobenius error norm is compared well with the optimal bound.
-The mere existence of such LRA algorithms is surprising for $n\times n$ SPSD matrices, whose class includes matrices $A^{*}A$, defined by nonsingular $n\times n$ matrices $A$ with $n^{2}$ independent entries.
-Notice that the superfast algorithm of [15, Part III] for SPSD matrices is deterministic, but its spectral and Frobenius output error norms exceed their optimal values by at most a factor of $n$.
-Unfortunately, one cannot compute LRA superfast for every matrix: already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [26, 27, 28]).
-If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0.
-1.2 Our LRA algorithm
+The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [10, Secs. 10 and 11], [30, Ch. 6.2], [5, Thm. 3.4]).
+1.3 Acceleration of randomized LRA by using sparse sketches
 Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$.
-We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$ of size $m\times n$ such that $r<\rho\ll\sqrt{n}$.
-Our algorithm is superfast and, for a large subclass of $(r,\rho)$-matrices $M$, its output very low rank approximation (VLRA) is near-optimal under spectral norm.
-We rely on combining three surprisingly simple observations, namely, that (i) random sampling with oversampling is the only stage where HMT, TYUC, and GN algorithms are not superfast for VLRA; (ii) we can perform this stage superfast by using sparse test matrices (in particular abridged SRHT matrices of Appendix D), and (iii) while the accuracy of the VLRA output in step (ii) tends to be unsatisfactory, we can readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$.
+We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$.
+We rely on combining three simple observations, namely, that (i) random sampling with oversampling is the bottleneck stage of HMT, TYUC, and GN algorithms, (ii) which we can greatly accelerate by using sparse sketch matrices (in particular abridged SRHT matrices of Appendix C), and (iii) while the accuracy of the LRA output in step (ii) tends to deteriorate with sparse sketch matrices, we readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$.
 Our LRA Alg.
@@
 2 relies on these observations.
-We compute at first (1) a crude but reasonably close rank-$\rho$ approximation $M(\rho)$ of $M$ and then (2) the rank-$r$ truncation $M(\rho)_{r}$ of $M(\rho)$, which we output as rank-$r$ approximation of $M$.
-LRA via random sampling with oversampling fits this outline, but based on above observations (i)–(iii) we run the algorithm superfast for $\rho\ll\sqrt{n}$ and argue that its output VLRA is near-optimal for a large subclass of $(r,\rho)$-matrices.
-We can perform stage (1) of the algorithm superfast by applying any of the HMT, TYUC and GN algorithms for $k:=\rho\ll\sqrt{n}$ with sparse test matrices such that their multiplication by $M$ is superfast (see [13], [33, Sec.
-3.9], and our Sec.
-3).
-Apart from the stage of multiplication by test matrices, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$, and obtain Moore-Penrose pseudoinverses of a $k\times\ell$ matrices ${\bf\Psi Q}$ in [33] and $R_{\epsilon}$ in [20] by using $O(k^{2}\ell)$ flops (see [7, Fig.
-8.6.1]).
-These computations are superfast where $(k+p)^{2}\ll n$ and $k^{2}\ll n$, respectively, for $\ell\leq 2k$, say.
-At Stage (2) we apply Alg.
-B.1.
-It computes SVD and $r$-truncation of $M(\rho)$ by using about $(m+n)\rho$ scalars and $O(mr^{2})$ flops, that is, $\ll m\sqrt{n}$ scalars and $\ll mn$ flops for $\rho\ll\sqrt{n}$.
+It first computes (1) a crude rank-$\rho$ approximation $M(\rho)$ of $M$; then (2) $r$-top SVD of $M(\rho)$, output as rank-$r$ approximation of $M$.
+LRA via random sampling with oversampling fits this outline.
+Apart from the stage of multiplication by sketch matrices, performed at a dominated cost, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$.
+Unlike the number of scalars, the bound on the number of flops decreases greatly where $k+p\ll n$ and $k\leq\ell\ll n$ for HMT and for TYUC and GN algorithms, respectively.
+Here and hereafter $a\ll b$ means that $b$ exceeds $a$ greatly in context, e.g., $a=o(b)$ as $b\rightarrow\infty$.
+In particular the algorithms HMT and TYUC/GN run at sublinear cost, that is, use much fewer than $mn$ scalars and flops, where we seek Very Low Rank Approximation (VLRA), namely, where $k+p\ll\sqrt{n}$ or $kl\ll\sqrt{n}$, respectively.
+Alg.
+B.1 of Appendix B performs Stage (2) by using about as many scalars and $O(m\ell^{2})$ flops.
+We can decrease this bound to $O(mk\ell)$ (which is sublinear for $kl\ll\sqrt{n}$) by applying TUYC or GN algorithms to compute near-optimal rank-$r$ approximation of $M(\rho)$ rather than its $r$-top SVD at stage (2) of Alg.
+2.1.
+1.4 Accuracy of the output LRA
 By virtue of Thm. 2.3, the spectral norm $||\cdot||:=||\cdot||_{2}$ of the output error matrix of Alg. 2.1 satisfies
@@
 2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$.
-For $(r,\rho)$-matrices, this factor can be large only if the superfast computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails.
-It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse test matrices tend to be unsatisfactory as LRA [13], [33, Sec.
+For $(r,\rho)$-matrices, this factor can be large only if the computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails.
+It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse sketch matrices tend to be unsatisfactory as LRA [11], [27, Sec.
 3.9], but in our numerical tests in Sec.
 5 accuracy of the output matrices of stage (1) of Alg.
-2.1 tended to be raised to near-optimal level at stage (2).
+2.1 tended to be raised at stage (2) to near-optimal level for a large sub-class of $(r,\rho)$-matrices $M$.
 1.5 Iterative LRA
 In Sec.
-4 we apply our LRA algorithm to the output error matrix, hoping to decrease its norm if it is too large.
+4 we apply our LRA algorithm (or we can apply any efficient LRA algorithm instead) to the output error matrix of a crude LRA computed at stage (1), to decrease its norm.
 Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs.
-The algorithm is essentially reduced to recursive applications of Alg.
+We reduce the algorithm essentially to recursive applications of Alg.
 2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance.
@@
 4 and 5.
-We devised this algorithm to complement and strengthen Alg.
-2.1, but it is essentially reduced to recursive applications of Alg.
-2.1.
-Our upper estimates for the accuracy of both algorithms are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
+Our upper estimates for the accuracy of both Alg. 2.1 and iterative refinement are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
 In our tests, the output accuracy of our iterative LRA applied to $(r,\rho)$-matrices improved significantly at the second iteration, little improved afterwards, and overall error norm was never less than with Alg.
@@
 1.6 Related works
-Our study was prompted by extensive work on LRA in the last decades (see [11, 33, 18] and the bibliography therein) and in particular by fast LRA based on random sampling and oversampling.
-To accelerate these algorithms we use abridged SRHT matrices studied in the papers [26, 27, 28], which in turn extends the earlier study of Gaussian elimination without pivoting in [29, 31].
-Our iterative LRA extends the algorithms of [16, 25], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [32, Secs.
-3.3.4 and 4.2.5], [10, Ch.
+Our iterative LRA extends the algorithms of [13, 19], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [26, Secs.
+3.3.4 and 4.2.5], [9, Ch.
 12 and Sec.
-20.5], [7, Secs.
-3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [32, page 223 and the references on page 225]).
+20.5], [6, Secs.
+3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [26, page 223 and the references on page 225]).
 Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [18, 24].
-Actually, we first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.
+Such compression is commonly used for LRA with oversampling.
+For sparse test matrices we used abridged SRHT matrices studied in the papers [20, 21, 22], which in turn extends the earlier study of Gaussian elimination without pivoting in [23, 25].
+Unlike our presentation in Secs.
+1.3 and 1.5 as well as in Secs.
+2 and 4, we actually first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.
 1.7 Organization of our paper
 We present our accelerated oversampling LRA algorithm in Sec.
-2, discuss superfast computation of a crude LRA in Sec.
+2, discuss fast computation of a crude LRA in Sec.
 3, present our iterative LRA algorithm in Sec.
@@
 6 to conclusions.
-In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails.
-In Appendix B we compute $r$-top SVD of an LRA of a low-rank matrix superfast.
-In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA.
-In Appendix D we generate abridged SRHT matrices, whose incorporation as test matrices makes random sampling LRA superfast.
+In Appendix A we describe some small families of matrices for which any LRA algorithm fails unless it access all entries of an input matrix.
+In Appendix B we fast compute $r$-top SVD of an LRA of a low-rank matrix.
+In Appendix C we generate abridged SRHT matrices, whose incorporation as sketch matrices makes random sampling LRA run faster.
 2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
@@
 Theorem 2.2.
-[7, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
+[6, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
 $|\sigma_{j}(M+E)-\sigma_{j}(M)|\leq||E||~{}{\rm for}~{}j=1,\dots,\min\{m,n\}.$
-In the following algorithm we elaborate upon the outline of Sec. 1.2 assuming that we are given two black box subroutines that output a crude LRA of a matrix.
+In the following algorithm we elaborate upon the outline of Sec. 1.3 assuming that we are given black box subroutines that output a crude LRA and an estimate for a spectral norm of a matrix.
 Algorithm 2.1.
@@
   2. 2.
-    Compute its $r$-top SVD (cf. Appendix B), which is SVD of $X:=M(\rho)_{r}$.
+    Compute its $r$-top SVD (cf. [19, Appendix B]), which is SVD of $X:=M(\rho)_{r}$.
   3. 3.
@@
 The error norm $||E||=||M-X||$ (cf. Eqn. (2.2)) satisfies
-$||E||\leq\sigma_{r+1}(M)+2\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||,$ (2.3)
+$||E||\leq\sigma_{r+1}(M)+2\nu~{}{\rm where}~{}\nu=\nu(M,\rho):=||M-M(\rho)||,$ (2.3)
 which turns into (1.3) for $\sigma_{r+1}(M)=||M-M_{r}||$.
@@
 Remark 2.1.
-Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$, but can be very pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
+Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$ but can be quite pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
 Remark 2.2.
-[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should keep $\rho$ small enough to perform at stage 2 fast or superfast.
+[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should choose it smaller to perform stage 2 faster.
 For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
@@
 2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$.
-In practice, the choice of integers $\rho$ and then $r$ should adaptive, towards minimizing the error norms of the resulting rank-$\rho$ and rank-$r$ approximations of $M$.
-3 LRA by means of random sampling with sparse multipliers
-The random sampling LRA algorithms of [11, 33] are close to optimal with a high probability but are not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT.
-In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations to superfast level by using sparse test matrices, e.g., Ultra-Sparse Rademacher test matrices [33, Sec.
-3.9], [3], which depend on a positive integer parameter $s$, such that a test matrices are sparse and support superfast LRA for smaller values $s$.
-It is a challenge to support our VLRA with these algorithms and test matrices, but in this paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix D and also supporting superfast VLRA.
-For other candidate sparse test matrices see [11, Remark 4.6], [7, Sec.
+In practice, the choice of integers $\rho$ and then $r$ should be adapted towards minimizing the error norms of the computed rank-$\rho$ and rank-$r$ approximations of $M$.
+3 LRA by means of random sampling with sparse sketch matrices
+The random sampling LRA algorithms of [10, 27] are close to optimal with a high probability under various customary choices of dense sketch matrices such as Gaussian, Rademacher, SRHT, and SRFT.
+In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations by using sparse sketch matrices, such as Ultra-Sparse Rademacher sketch matrices [27, Sec.
+3.9], [2], which depend on a positive integer parameter $s$, such that sketch matrices are sparse and support faster LRA for smaller values $s$.
+In the present paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix C and also supporting faster VLRA.
+For other candidate sparse sketch matrices see [10, Remark 4.6], [6, Sec.
 5.1.8].
-According to [13], such acceleration tends to make the output LRAs somewhat less reliable, and indeed in our tests LRAs computed at stage 1 of Alg.
-2.1 tended to be cruder with abridged SRHT test matrices than with Gaussian random ones.
+According to [11], such acceleration tends to make the output LRAs somewhat less reliable, and in our tests LRAs computed at stage 1 of Alg.
+2.1 tended to be cruder with abridged SRHT sketch matrices than with Gaussian random ones.
 Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
@@
 Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec.
-1.3.
-We allow at most $k$ refinement steps for a fixed tolerance $k$.
+1.5.
+As a basis we apply Alg.
+2.1, but other efficient LRA algorithms can be applied instead of it.
+We allow at most $h$ refinement steps for a fixed tolerance $h$.
 As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
@@
 2.1, except that the input of Alg.
-4.1 additionally includes a positive integer $k$, which bounds the number of recursive steps allowed, and an Alg.
+4.1 additionally includes a positive integer $h$, which bounds the number of recursive steps allowed, and an Alg.
 $\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
@@
   3. 3.
-    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This compresses $M_{i}(\rho)$.]
+    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This enables us to compress $M_{i}(\rho)$.]
   4. 4.
     Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if it satisfies the tolerance bound, that is, if the $\mu_{i}:=\mu_{i}(M,\rho)\leq\mu$.
-    Otherwise output FAILURE if $i=k$.
+    Otherwise output FAILURE if $i=h$.
   5. 5.
@@
 5.3 we test a variant of Alg.
-4.1 where we fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
+4.1 where we fix $h=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
 In the following extension of Thm. 2.3 we estimate the output error norm of Alg. 4.1.
 Theorem 4.1.
-For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,k-1.$ Then it holds that
+For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,h-1.$ Then it holds that
 $\mu_{i+1}:=||E_{i+1}||\leq\sigma_{r+1}(M)+2\nu_{i}.$ (4.2)
@@
 The singular values $\sigma_{\rho+1}(E_{i})$ satisfy
-$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,k-1.$ (4.3)
+$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,h-1.$ (4.3)
 Remark 4.3.
@@
 […]
-Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [11, Sec. 7.1] for more details).
+Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [10, Sec. 7.1] for more details).
 Fig. 1 shows the distribution of the top 50 singular values of these three matrices.
@@
 Figure 1: Spectra of singular values of some real-world matrices
-We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [34, Sec. 7.3.2].
+We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [28, Sec. 7.3.2].
 The $10738\times 5001$ matrix StreamVel is involved in numerical simulation (DNS) on a coarse mesh of the 2D Navier–Stokes equations for a low-Reynolds number flow around a cylinder.
@@
 By letting $\Sigma=\textrm{diag}(u)$, $u_{i}=1$ for $i=1,2,\dots,20$, and $u_{i}=\frac{1}{(1+i-20)^{2}}$ for $i>20$, we arrived at the matrices with slowly decaying spectra.
-For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [34, Sec.
+For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [28, Sec.
 7.3.1], which involved effective rank parameter $R=20$.
@@
 In our tests of Algs. 2.1 and 4.1 we used the following two options at their stage 1 for computations of a crude rank-$\rho$ approximation of a matrix $M$.
-(i) Fast (although not superfast) randomized algorithm of [33] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
-(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{C}^{(2\rho)\times 1024}$ and $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
+(i) Fast randomized algorithm of [27] with Gaussian random sketch matrices $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
+(ii) Its accelerated randomized variant where abridged (length 3) SRHT sketch matrices $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
 5.3 Test results for Algs. 2.1 and 4.1
@@
 For Alg. 2.1 we also display the standard deviation.
-For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg.
+For both algorithms we repeated the tests 100 times per test set with independent random choices of sketch matrices $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg.
 2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg.
@@
 We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
-In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT test matrices, which were more or less in the same range for both matrix types.
-All our tests with the synthetic input matrices from [34, Sec.
-7.3.1] wit abridged SRHT test matrices have failed, that is, have output large error norms.
-In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random test matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
+In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT sketch matrices, which were more or less in the same range for both matrix types.
+All our tests with the synthetic input matrices from [28, Sec.
+7.3.1] wit abridged SRHT sketch matrices have failed, that is, have output large error norms.
+In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random sketch matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
 Alg.
-2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT test matrices.
+2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT sketch matrices.
 This could have been predicted because the spectrum of the singular values $\sigma_{j}(M)$ of these matrices $M$ is flat for $j>20$ (decreasing roughly from $10^{-13}$ to between $10^{-15}$ and $10^{-16}$); in Table 5.1 we display the test results for the matrix Shaw and target rank $r=19$, which are about as favorable as for Gravity and Slow Decay.
@@
 This was not the case just for the matrices with flat spectra such as the matrices Shaw for $r=20$, High Noise, and Exp Decay Slow.
-Input Matrix Multiplier $\rho$
+Input Matrix Sketch Matrix $\rho$
 $2r$ $3r$ $4r$ $5r$
@@
 Table 5.1: Test results for Alg. 2.1 on Gravity, discretized single layer potential operator (SLP), Shaw, and synthetic matrices Fast Decay and Slow Decay, with fast and slowly decaying spectra.
-Input Matrix Multiplier 1st Itr. 2nd Itr. 3rd Itr.
+Input Matrix Sketch Matrix 1st Itr. 2nd Itr. 3rd Itr.
 Fast Decay Abr. SRHT 3.155000 1.0000 1.0000
@@
 Exp Decay Fast 1.3121 $\pm$ 1.4989E-01 1.$0$$\pm$ 6.4663E-11 1.$0$$\pm$ 3.6020E-16 1.$0$$\pm$ 3.0986E-16
-Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian test matrices.
+Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [28, Sec. 7.3.1] using Gaussian sketch matrices.
 Input Matrix 1st Itr. 2nd Itr. 3rd Itr.
@@
 Exp Decay Fast 1.4235 1.0000 1.0000
-Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian multipliers.
-Input Multiplier $\rho$
+Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [28, Sec. 7.3.1] using Gaussian sketch matrices.
+Input Sketch Matrix $\rho$
 Matrix Type $2r$ $3r$ $4r$ $5r$
@@
 Gaussian 1.0026 $\pm$ 1.3856E-03 1.0005 $\pm$ 2.1129E-04 1.0002 $\pm$ 6.6260E-05 1.0001 $\pm$ 2.4207E-05
-Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [34, Sec. 7.3.2]
-Input Matrix Multiplier Type 1st Itr. 2nd Itr. 3rd Itr.
+Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [28, Sec. 7.3.2]
+Input Matrix Sketch Matrix Type 1st Itr. 2nd Itr. 3rd Itr.
 StreamVel Abr. SRHT 2.9700 1.0739 1.0300
@@
 Gaussian 2.6202E+01 1.0302 1.0176
-Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [34, Sec. 7.3.2].
+Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [28, Sec. 7.3.2].
 6 Conclusions
-Rank-$r$ approximation of an $m\times n$ matrix $M$ is only interesting if there is a gap in the spectrum of its top (largest) $r+1$ singular values $\sigma_{j}(M)$.
-For the $(r,\rho)$-matrices $M$ having also a gap in the next segment of their spectra – from $(r+1)$st to $(\rho+1)$st largest singular values – randomized oversampling algorithms output quite accurate LRA.
-We extended these observations towards devising superfast (sublinear cost) variation of LRA algorithms based on random sampling and oversampling with sparse test matrices and to improving accuracy of the resulting LRA in the case of $(r,\rho)$-matrices $M$ where $r<\rho\ll\sqrt{n}$.
-We provided some formal and experimental support for the efficiency of the resulting LRA algorithms.
 There are various natural extensions of our study, which can be practically competitive.
-E.g., one can devise and test similar variations of the TYUC and GN algorithms with sparse test matrices such as Ultra-Sparse Rademacher test matrices [33, Sec.
-3.9] or sparsified random Givens rotations by Rokhlin and Tygert (see [11, Remark 4.6]).666[26, 27, 28] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes.
-Our study in Secs. 4 and 5 has showed no benefits for using iterative refinement of LRA versus oversampling so far, but one can keep comparing their efficiency.
+E.g., one can devise and test combinations of Alg.
+2.1 and of the TYUC and GN algorithms with various sparse sketch matrices such as Ultra-Sparse Rademacher sketch matrices [27, Sec.
+3.9], sparsified random Givens rotations by Rokhlin and Tygert (see [10, Remark 4.6]), or even further sparsified count sketch matrices in Clarkson and Woodruff [2, 3].444[20, 21, 22] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes.
+Instead of random sampling LRA with sparse sketch matrices one can apply Cross-Approximation algorithm for LRA (see [17] and the references therein).
+Our comparative study has showed no significant benefits of applying iterative refinement to LRA versus standalone Alg. 2.1, but one can further explore this comparison.
 Appendix
-Appendix A Small families of matrices that are hard for superfast LRA
-Any superfast algorithm fails to compute a close LRA of the following small families of matrices.
+Appendix A Small families of matrices that are hard for fast LRA
+Any algorithm that does not access all entries of an input matrix fails to compute a close LRA of the following small families of matrices.
 Example A.1.
@@
 We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family.
-Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$.
-Therefore it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2.
-Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank.
-Likewise, we can verify that any superfast randomized LRA algorithms fail fails on those matrix families with error probability not close to 0.
+Now fix any algorithm that does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$.
+Such n algorithm outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2.
+We arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank.
+Likewise, we can verify that any randomized LRA algorithms that does not access ail entries of an input matrix fails on those matrix families with error probability not close to 0.
 Remark A.1.
@@
 The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf.
-[7, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$.
+[6, Figure 8.6.1]), running particularly fast where $k\ll n\leq m$.
 Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$.
 One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
-Appendix C CUR decomposition, CUR LRA, and its link to top SVD
-C.1 CUR decomposition and CUR approximation
-Next we briefly revisit CUR LRA, which is particularly memory efficient.
-For five integers $k$, $\ell$, $m$, $n$, and $r$ satisfying (1.2), an $m\times n$ matrix $M$, and two sets $\mathcal{I}\subseteq\{1,\dots,m\}$ and $\mathcal{J}\subseteq\{1,\dots,n\}$ of cardinalities $k$ and $\ell$, respectively, define the submatrices
-$R:=M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},C:=M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{}{\rm and}~{}G:=M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}$
-of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$, although we recall notation $G$ for nucleus and CGR for the factorization in [21], originated with the decomposition itself in the 1990s [8].
-In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
-Given a $k\times\ell$ generator submatrix $G$ of $M$, we can compute CUR approximation of $M$ by using $k\ell$ scalars and $O(k\ell\min\{k,\ell\})$ flops.
-Theorem C.1.
-(See [28].) For matrices $C$, $U$, and $R$ defined above, $M=CUR$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
-If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large.
-C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
-The algorithm of Appendix B transforms a CUR decomposition into a top SVD.
-The next algorithm makes converse transition.
-It uses $(m+n+r)r$ scalars and order of $(m+n)r^{2}$ flops to compute CUR decomposition having an $r\times r$ generator for a matrix given with its $r$-top SVD, for $r\leq n\leq m$.
-By combining this algorithm with that of Appendix B we can transform an LRA of a matrix into its CUR LRA at a low computational cost.
-Algorithm C.1.
-[From $r$-top SVD to CNR decomposition.] [^7] We denote nucleus by $N$ rather than $U$ or $G$ to avoid conflicts with the factor $U$ in SVD and generator $G$ of CUR of Sec. C.1.
-Input:
-Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{C}^{m\times n}$, $\Sigma\in\mathbb{C}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{C}^{m\times r}$, and $V\in\mathbb{C}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
-Output:
-Three matrices $C\in\mathbb{C}^{m\times r}$, $N\in\mathbb{C}^{r\times r}$, and $R\in\mathbb{C}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
-$M=CNR.$
-Computations:
-By applying to the matrices $U$ and $V$ the algorithms of [6] or [23] compute two nonsingular $r\times r$ submatrices $U_{\mathcal{I},:}$ and $V^{*}_{:,\mathcal{J}}$.
-Compute and output the CUR factors $C=U\Sigma V^{*}_{:,\mathcal{J}}$, $R=U_{\mathcal{I},:}\Sigma V^{*}$, and $N:=(V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1}$.
-Correctness verification.
-Substitute the expressions for $C$, $N$ and $R$ and obtain $CNR=(U\Sigma V^{*}_{:,\mathcal{J}})((V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1})(U_{\mathcal{I},:}\Sigma V^{*})=U\Sigma V^{*}$.
-Cost bounds. The algorithm uses $O(mr^{2})$ flops and $(m+n+r)r$ scalars; these cost bounds are sublinear where $r^{2}\ll n$ and are dominated at stage 2.
-Let us also estimate the norm of the nucleus $||N||$.
-Observe that $\operatorname{rank}(V_{:,\mathcal{J}})=\operatorname{rank}(\Sigma)=\operatorname{rank}(U_{\mathcal{I},:})=r$ and $||N||\leq||V_{:,\mathcal{J}}^{-1}||~{}||\Sigma^{-1}||~{}||U_{\mathcal{I},:}^{-1}||$ and recall that $||\Sigma^{-1}||=||M^{+}||=1/\sigma_{r}(M)$.
-Write $t_{q,s,h}^{2}:=(q-s)sh^{2}+1$, allow any choice of $h>1$, say, $h=1.1$, and then recall that $||U_{\mathcal{I},:}^{+}||\leq t_{m,k,h}^{a}$, $||(V_{:,\mathcal{J}}^{*})^{+}||\leq t_{n,\ell,h}^{a}$, and consequently
-$||N||\leq t_{m,r,h}^{a}t_{n,r,h}^{a}/\sigma_{r}(M)$
-where $a=1$ if we apply the algorithms of [6] at stage and $a=2$ if we apply those of [23].
-Appendix D Generation of abridged SRHT matrices
+Appendix C Generation of abridged SRHT matrices
 In this section we specify the family of abridged SRHT matrices, used in our tests for Algs.
@@
 They are defined by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf.
-[17, Sec.
+[14, Sec.
 3.1]).
@@
 Namely, we write $H_{d,0}:=I_{n/2^{d}}$ and then specify the following recursive process:
-$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (D.1)
+$H_{d,0}:=I_{n/2^{d}},~{}H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{}{\rm for}~{}i=0,1,\dots,d-1,$ (C.1)
 For any fixed pair of $d$ and $i$, each of the matrices $H_{d,i}$ is orthogonal up to scaling and has $2^{d}$ nonzero entries in every row and column; we can compute the product $MH$ for an $n\times k$ submatrix $H$ of $H_{d,d}$ by using less than $km2^{d}$ additions and subtractions.
@@
 The first $k$ columns of $H_{d,d}$ for $r\leq k\leq n$ form a $d$-Abridged Subsampled Randomized Hadamard Transform (SRHT) matrix $H$, which turns into a SRHT matrix for $d=t$, where $k=r+p$, $r$ is a target rank and $p$ is the oversampling parameter (cf.
-[11, Sec.
+[10, Sec.
 11]).